### v2

# I Learn Better If You Speak My Language: Understanding the Superior Performance of Fine-Tuning Large Language Models with LLM-Generated Responses

Abstract
This paper explores an intriguing observation: fine-tuning a large language model (LLM) with responses generated by a LLM often yields better results than using responses generated by humans. We conduct an in-depth investigation to understand why this occurs. Contrary to the common belief that these instances is simply due to the more detailed nature of LLM-generated content, our study identifies another contributing factor: an LLM is inherently more "familiar" with LLM generated responses. This familiarity is evidenced by lower perplexity before fine-tuning. We design a series of experiments to understand the impact of the "familiarity" and our conclusion reveals that this "familiarity" significantly impacts learning performance. Training with LLM-generated responses not only enhances performance but also helps maintain the model’s capabilities in other tasks after fine-tuning on a specific task.
I Learn Better If You Speak My Language: Understanding the Superior Performance of Fine-Tuning Large Language Models with LLM-Generated Responses
Xuan Ren xuan.ren@adelaide.edu.au                        Biao Wu biaowu165534@gmail.com                        Lingqiao Liuthanks: Corresponding author. lingqiao.liu@adelaide.edu.au
1 Introduction
Recent research has demonstrated that a large language model (LLM) can generate training data for another LLM Dai et al. (2023); Edwards et al. (2022); Møller et al. (2024); Guo et al. (2023); Ubani et al. (2023); Agrawal et al. (2023); Xu et al. (2023); Kieser et al. (2023); Taori et al. (2023); Peng et al. (2023); Xu et al. (2024). This approach offers a method for transferring knowledge from a larger model to a smaller one, or for creating supplementary training materials, such as rationales Zhang et al. (2024); Kang et al. (2023); Li et al. (2022), or refined step-by-step reasoning Hsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023).
The quality of data generated by large language models (LLMs) is considered to be very high and conducive to effective trainingLee et al. (2023). The prevailing view often attributes the successful training results of LLM-generated data to the fact that LLMs provide more details than human-annotated data, such as ’chain-of-thought-style’ reasoning processHsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023). However, our experiments have found numerous counterexamples. As shown in Figure 1, target labels that include more details and utilize chain of thought styles do not necessarily result in better training outcomes. Through experiments on various tasks, we have observed that when LLMs directly generate target labels, the training results are often superior to those obtained by transforming ground truth into more detailed, step-by-step training. This indicates that whether the answer generation is executed in the LLM’s own style or is based on a rewritten human style is a key factor affecting training outcomes.
Figure 1: This figure illustrates the training outcomes associated with various data generation methods. As demonstrated by the examples shown, the inclusion of additional details does not necessarily translate into enhanced performance.
This observation motivates us to investigate the underlying reason for this phenomenon. We found that whether the LLM is familiar with the target label is a crucial factor influencing model training.
Our main hypothesis are as follows:
  • Given the same question, an LLM is inherently familiar with responses generated by itself or other LLMs.
  • an LLM can better adapt to and leverage responses in its familiar styles, leading to improved outcomes.
The following section of this paper will examine the hypothesis through a series of experiments. We will measure ’familiarity,’ conduct ablation studies by creating different variants of LLM-generated responses, and evaluate their performance to understand various factors in the process. This will provide insights into our hypothesis.
2 Empirical Study Protocols
The purpose of this paper is to understand and investigate the aforementioned phenomenon rather than to propose a new method. Our study’s methodology is based on empirical analysis. We begin by introducing our study protocols, including the datasets and evaluation approaches used.
2.1 Evaluation approaches
Without loss of generality, we denote the training data as question and response pairs $(q,r)$. We will use a subset of the LLaMA2 benchmark Touvron et al. (2023) to create the training dataset. We will use a subset of the LLaMA2 benchmark [34] that contains human-labeled explicit reasoning processes to create the training dataset. The question and response pairs in these datasets are human-constructed. Additionally, we will use an LLM to generate new responses, following $r^{\prime}=LLM(q,P)$, where $LLM$ denotes the language model and $P$ denotes the prompt used to create the response. We will explore different $P$ in the following experiment. This creates a synthetic version of dataset $\{(q,r^{\prime})\}$.
Our training approach for the LLaMA2-13b-chatTouvron et al. (2023) and Mistral-7b-Instruct-v2Jiang et al. (2023) models involves specific adaptations to optimize performance. Notably, the warm-up period is set to 10% of the total training steps consistently across our experiments. For the LLaMA-13b-chat model training on the HumanEval dataset, we set the learning rate to 2e-4. This adjustment was necessary as lower rates proved insufficient for the model to effectively learn from this dataset. Given the small size of the training datasets, such as MBPP and HumanEval, we reduced the batch size to 10 to accommodate the limited data. In contrast, for general fine-tuning of the LLaMA2-13b-chat and Mistral-7b-Instruct-v2 models, we employ a default learning rate of 2e-5, a batch size of 32, and a warm-up period comprising 10% of the total steps. Additionally, we use a cosine learning rate schedule and fine-tune only the Q and V matrices of the LORAHu et al. (2021) parameters with a rank of 8. All models are trained and evaluated using half-precision.
We evaluate the fine-tuned model in two different ways. One is called in-domain evaluation, which assesses performance on the tasks of the training set. The other is called cross-task evaluation, which evaluates performance across tasks on other datasets. The purpose of using cross-domain evaluation is to test if the model can benefit from the training process of a similar task and if the model will forget its capability on other tasks after training on a specific task.
3 Experiment setup
3.1 Dataset and Evaluation
We utilized the LLaMA2 benchmark Touvron et al. (2023) to select five tasks that contain human-labeled explicit reasoning. These were evaluated based on accuracy to assess the effectiveness of different data construction methods. The term ’dataset with explicit reasoning’ refers to datasets that include a detailed reasoning process in their annotations. This category includes GSM8KCobbe et al. (2021), MATHHendrycks et al. (2021), HumanEvalChen et al. (2021), MBPPAustin et al. (2021), and ECQA Aggarwal et al. (2021), where the ground truth comprises a complete human-labeled problem-solving process or rationale, not just the final answer. For example, the ground truth in GSM8K involves comprehensive mathematical reasoning, and the coding datasets’ ground truths consist of complete codes, which are logical sequences by nature. We included coding datasets like MBPP and HumanEval in this category, as developing code inherently involves logical reasoning, even though it may not always be presented in the form of a traditional chain of thought. Our model was trained on one dataset and subjected to in-domain testing on the same dataset, as well as cross-domain testing on other datasets. Our selection includes all coding and mathematics datasets within the LLaMA2 benchmark. We chose to use the ECQA dataset because, unlike the CQATalmor et al. (2018) dataset included in the LLaMA2 benchmark, ECQA includes rationales provided by human annotators.
Code: We evaluated coding ability using the HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) datasets. We report the pass@1 accuracy for each dataset. Note that the HumanEval dataset contains only 164 testing data points with no training data. We divided the training task into two parts: in the first part, we trained on the first 82 data points and tested on the last 82. In the second part, we reversed this order. Ultimately, we averaged the test results from both parts. For the MBPP-full dataset, as opposed to the MBPP-sanitized version which includes fewer training data points, we utilized all available training data (374 data points) and testing data (500 data points)
Commonsense Reasoning: For Commonsense QA task, we selected the ECQA Aggarwal et al. (2021) dataset, which provides explanations for the answers from CQATalmor et al. (2018) dataset. We combined these explanations with the multiple-choice answers to form the ground truth training data. We used the first 1000 training and test data points for training and evaluation.
Math: We conducted training and evaluations on the GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021) datasets. The MATH dataset encompasses several categories. We conducted training and evaluations exclusively within the algebra category. For the evaluation, we developed a script that compares the ground-truth answer with the model’s final prediction. This script accurately evaluates responses that are numerical. However, in the MATH dataset, some answers contain inequalities or complex mathematical expressions. We filtered out these data points and used only those with numerical answers for both training and evaluation. After excluding data points with non-numerical answers, we used the first 1000 training data points for training and all the testing data points for testing. As a result, we obtained 1000 training and 1314 testing data points for GSM8K, and 1000 training and 752 testing data points for MATH (algebra).
Evaluation Script:
To facilitate automatic evaluation across all models and datasets, we developed a specialized prompt format requiring models to conclude their predictions with ’Final Answer: answer.’ For instance, the output should read ’Final Answer: 100’ for an answer of 100. Our evaluation script extracts the text following the ’Final Answer:’ substring. For the ECQA task, it verifies if the first text after ’Final Answer:’ correctly matches one of the answer options. For mathematical problems including the GSM8K and MATH datasets, the script checks if the number after the substring ’Final Answer:’ aligns with the ground truth number. We utilize the built-in testing script for the HumanEval dataset to ensure result accuracy and employ our custom scripts for all other tasks. By adapting the training data’s question and answer formats to this style, our models maintain consistent response formatting across different tasks. The evaluation scripts for HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) require answers in a specific format. Models that have not been trained on this format struggle to follow it correctly, even when provided with in-context examples. Consequently, we do not conduct cross-domain evaluations on the MBPP and HumanEval datasets.
4 Observations and Hypothesis:
4.1 Observation 1: Superior Performance of GPT-4 Over Human-Annotated Labels
Method Training Dataset and Model Type GSM8K Math Algebra ECQA HumanEval MBPP
Groundtruth GSM8K, Mistral 0.434 0.162 0.594
GPT-4 Answer Directly 0.597 0.246 0.597
Groundtruth GSM8K, Llama 0.364 0.141 0.565
GPT-4 Answer Directly 0.428 0.128 0.575
Groundtruth Math algebra, Mistral 0.264 0.206 0.554
GPT-4 Answer Directly 0.553 0.302 0.608
Groundtruth Math algebra, Llama 0.36 0.126 0.575
GPT-4 Answer Directly 0.35 0.150 0.561
Groundtruth ECQA, Mistral 0.258 0.134 0.68
GPT-4 Answer Directly 0.462 0.223 0.722
Groundtruth ECQA, Llama 0.132 0.0798 0.631
GPT-4 Answer Directly 0.379 0.156 0.656
Groundtruth HumanEval, Mistral 0.3625 0.191 0.583 0.323
GPT-4 Answer Directly 0.313 0.163 0.581 0.366
Groundtruth HumanEval, Llama 0.0705 0.083 0.528 0.146
GPT-4 Answer Directly 0.1245 0.105 0.552 0.159
Groundtruth MBPP, Mistral 0.392 0.176 0.527 0.276
GPT-4 Answer Directly 0.399 0.186 0.568 0.354
Groundtruth MBPP, Llama 0.328 0.132 0.556 0.2
GPT-4 Answer Directly 0.351 0.140 0.574 0.202
Table 1: Comparison of human-annotated data versus data generated directly by GPT-4. We employ GPT-4 to generate responses using only the questions from the training datasets. For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation. In-domain performance is highlighted in grey. Data points significantly lower in accuracy by 15% or more relative to another method are marked in red.
Our study is inspired by a compelling observation: when a sophisticated LLM, such as GPT-4, is used to generate responses to questions, and these questions along with their generated responses are used as a synthetic dataset, fine-tuning smaller LLMs (we use both Llama2-13B-Chat and Mistral-7B-Instruct-v2) on this synthetic dataset often yields better performance than using the original dataset with human-provided answers. The detailed experimental results are presented in Table 1. As demonstrated, the majority of the low-performance data points (14 out of 15) marked in red reflect accuracies when training with human-annotated ground truth. Using synthetic datasets consistently yields higher in-domain performance compared to human-answered datasets. Notably, we observe a significant improvement in math-related tasks, often exceeding a 10% absolute increase in in-domain performance. Additionally, the model trained on the synthetic data achieves higher cross-domain performance. This observation motivates us to investigate the underlying reasons.
4.2 Observation 2: significantly lower perplexity of GPT-4 and other LLMs Over Human-Annotated Labels
Figure 2: Average Perplexity Comparison
We measure the perplexity of responses as assessed by the target (smaller) LLMs. An interesting observation, shown in Figure 2, is that, given the input questions from the responses generated by GPT-4 attain significantly lower perplexity compared to those provided by humans. To explore whether this phenomenon is unique to GPT-4 or common among LLM responses, we employed a zero-shot approach with the Mistral-7B-Instruct-v2 model and the LLaMA-13B-chat model. These models answered the same questions from the training set. As shown in Figure 2, their perplexity was also significantly lower than that of the human-labeled dataset. Lower perplexity indicates that LLMs are more familiar with responses generated by themselves or other LLMs than with human-generated responses.
4.3 Hypotheses
From Observation 2, we notice that given the same question, the average perplexity of responses generated by LLMs is consistently lower than the perplexity of ground truth annotated by humans. This observation leads us to pose our first hypothesis: When generating target labels for the same question, one LLM inherently more familiar with data generated by another LLM. We not only observed consistently lower perplexity from GPT-4 across tasks in Observation 2 but also noted that LLMs trained on GPT-4 generated responses consistently outperform the same LLMs trained using human-annotated data. This phenomenon leads us to propose our second hypothesis: An LLM performs better when trained on data it is familiar with.
5 Investigation of the proposed hypotheses
Method Training Dataset and Model Type GSM8K Math Algebra ECQA Perplexity token length
Groundtruth GSM8K, Mistral 0.434 0.162 0.594 4.54 128.88
GPT-4 Answer Directly 0.597 0.246 0.597 2.35 179.94
Step by step 0.574 0.224 0.591 2.27 282.9
Step by step transformation of GT 0.499 0.190 0.6 2.67 235.892
Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.88
GPT-4 Answer Directly 0.428 0.128 0.575 2.30 179.941
Step by step 0.402 0.134 0.553 2.19 290.92
Step by step transformation of GT 0.394 0.146 0.573 2.43 242.22
Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.83 163.36
GPT-4 Answer Directly 0.553 0.302 0.608 2.20 328.76
Step by step 0.533 0.239 0.583 2.249 451.99
Step by step transformation of GT 0.444 0.210 0.577 2.65 370.35
Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.15 163.36
GPT-4 Answer Directly 0.35 0.150 0.561 2.59 328.76
Step by step 0.344 0.145 0.593 2.55 450.08
Step by step transformation of GT 0.376 0.141 0.565 2.81 369.06
Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.3 73.51
GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.54
Step by step 0.481 0.203 0.71 4.66 325.15
Step by step transformation of GT 0.462 0.194 0.681 6.89 218.748
Groundtruth ECQA, Llama 0.132 0.0798 0.631 11.96 73.51
GPT-4 Answer Directly 0.379 0.156 0.656 5.33 176.54
Step by step 0.363 0.140 0.648 4.46 337.30
Step by step transformation of GT 0.312 0.108 0.628 5.55 226.99
Table 2: Performance comparison of models trained on data constructed by different methods across various tasks. Data points are labeled as low performance when their accuracy is more than 15% lower relative to the highest accuracy achieved in the same dataset using the same model.
In this section, we conduct an in-depth investigation about the aforementioned hypotheses. Our investigations are organized into two parts. In Section 5.1, we examine to what extend the familiarity contribute to higher performance of using LLM generated responses. Particularly, we conduct experiments to ablate the factor that LLM could introduce additional reasoning details to the original human-provided response. In Section 5.2, we experiment to determine if a model, when used alone without the assistance of a more advanced model, can rewrite ground truth into a form it is more familiar with, thereby enhancing training outcomes. The advantage of this experimental design is that it allows us to evaluate the effect of familiarity while removing any benefits that might arise from using a more advanced language model.
5.1 Understand the impact of “familiarity”
Experiment 1: Is observation 1 only rooted from more detailed reasoning steps provided by the LLM other than familiarity?
Our observation 1 suggests that using GPT-4 generated data leads to better training effectiveness than using human-annotated ground truth. Our second hypothesis posits that this enhanced learning by LLMs on LLMs responses could be attributed to familiarity with the data. It is a common belief that GPT-4 often includes more detailed content than human annotators, which could contribute to improved training outcomes. However, we question whether the superior results from GPT-4 are solely due to the addition of more details, or if familiarity plays a significant role as well.
To explore whether familiarity is a crucial factor, we designed the following experiment. We employed different methods to have GPT-4 generate data in a detailed, step-by-step manner. For each dataset, we established a control group: one group involved GPT-4 transforming ground truth into its own detailed style, while the other had GPT-4 mimic the style of the human-labeled ground truth, adding detail to build upon it in a step-by-step style. As shown in Table 2, in the vast majority of cases, the training effectiveness of the first group, which used GPT-4’s own style, was superior to that of the second group, which followed the human-annotated style. In addition, it is noteworthy that using the Mistral 7B model for training on math algebra, converting ground truth into step-by-step answers was even detrimental. Both of the step by step groundtruth transformation methods achieves accuracy below 24% on Math Algebra, much lower than 30.2% achieved by ’GPT-4 answer directly’(Gpt-4 answer directly represents employing GPT-4 to generate responses using only the questions from the training datasets. For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation.) We speculate that one possible reason for this is that adding ’more than enough details’ may sometimes complicate the model’s reasoning process. We have included additional experiments involving various GPT-4 generated variants in the appendixA, which generally conform to the patterns described above. For clarity and conciseness, Table 2 only includes 2 examples of the step-by-step variants.
After preliminary validation of our hypotheses, we made another intriguing observation: According to Table 2, the ’GPT-4 answer directly’ dataset, characterized by a shorter average token length of target labels compared to both ’step by step’ and ’step by step transformation of GT’ datasets, suggests fewer details in its content. Despite this, the performance of the data directly generated by GPT-4 frequently ranks among the best of all GPT-4 answer variants. Meanwhile, other data generation methods exhibited significantly lower performance than GPT-4’s direct answers on at least two tasks, as highlighted in red in Table 2. This suggests that, given a problem, the data directly generated by GPT-4 might yield the good training results. Artificially defined prompts that add information to the ground truth may lead to the generation of target labels that do not conform to the LLM’s linguistic or logical preferences.
Experiment 2: Direct measuring the impact of familiarity
Method Model Type GSM8K GSM8K Perplexity Math Algebra Math Algebra Perpleity ECQA ECQA Perplexity
GPT-4 answer with lower perplexity Mistral 0.597 2.35 0.301 2.20 0.722 6.35
GPT-4 answer with higher perplexity 0.547 5.58 0.231 4.66 0.702 11.13
GPT-4 answer with lower perplexity LLaMA 0.428 2.30 0.150 2.59 0.656 5.33
GPT-4 answer with higher perplexity 0.38 4.19 0.131 4.35 0.654 7.69
Table 3: Comparison between the GPT-4 answer with higher perplexity Vs. lower perplexity
This section introduces a method for assessing the influence of style ’familiarity’. We direct GPT-4 to produce several varient of responses to the same query, grouping them into two distinct datasets based on their perplexity levels—high and low. We then evaluate and compare the learning outcomes using these two sets. One concern of this experiment is that GPT-4 may produce answers with different level of details. To eliminate this factor, we first let GPT-4 to generate one answer and then explicitly direct it to paraphrase the generated answer. In this way, we can create responses that vary in perplexity but maintain the same semantic meaning.
After training on those two sets of training data, we examine their performance, shown in Table 3. As seen, training on the higher-perplexity training set consistently shows slightly lower performance than the lower perplexity counterpart, particularly noticeable in the in-domain performance on two math datasets. For example, in the algebra dataset from Mistral, the lower perplexity set achieves a 30% accuracy rate, while the higher perplexity set only reached 23%. Since they are only differed in the style of the text, this result demonstrated that the familiarity of LLMs with the stylistic aspects of target labels significantly affects training outcomes.
From these findings, we conclude that LLMs perform poorly when trained on unfamiliar data, explaining the suboptimal training results with human-annotated data.
Discussion: Why are LLMs often familiar with responses generated from other LLMs? The familiarity observed when training on data generated by GPT-4 might stems from the overlap in training data used during the pre-training stages of these models. Many LLMs, including GPT-4, share similar pre-training data, developing an understanding of specific output styles and content.
5.2 Achieving high familiarity by exploring the target LLM generated responses
Math Commensense Reasoning NLI
Method Model Type GSM8K Math Algebra ECQA ESNLI
Groundtruth Mistral 0.434 0.206 0.68 0.814
Correct predicted sample + groundtruth 0.491 0.227 0.578 0.557
correct preidcted sample only 0.449 0.257 0.586 0.698
Style Transfered Groundtruth 0.51 19.6 0.696 0.845
Table 4: Performance Comparison when not using GPT-4. For the style-transferred ground truth dataset, only 73.7% and 24% of the training data are used for training, as explained in Section5.2
In the experiments described above, we validated two hypotheses. However, all the studies mentioned use GPT-4 to generate responses, chosen for its advanced capabilities to directly produce correct answers and reliably follow diverse instructions. However, there is a concern regarding the "familiarity" factor: would this argument still hold without the involvement of an advanced LLM? If valid, the argument should apply to some extent even without using advanced LLMs.
This section seeks to develop a method that enhances the “familiarity” of responses without requiring more advanced large language models (LLMs) than the target LLM.
We included ESNLICamburu et al. (2018) as an additional dataset to better compare different data construction methods. The e-SNLI dataset enhances the SNLIBowman et al. (2015) dataset by adding human-annotated explanations to each instance. These annotations offer insights into the reasoning behind the labels of entailment, contradiction, or neutral, improving model interpretability and aiding in the development of more robust inference systems.
Use the same model for train and groundtruth style transfer We devised a method to rewrite the ground truth data in a style that resembles language model-generated responses using the Mistral-7B-Instruct-V2 model. The experiment results are shown in Table 4. Initially, we posed questions from the training dataset and recorded the answers that were correctly predicted by Mistral. We documented two such correct predictions. These two correct responses were then used as in-context examples to guide Mistral on how to rewrite ground truth data. Specifically, we used the two recorded predictions as example target labels and their corresponding questions as example questions, while the actual ground truth served as the example ground truth. The in-context prompt informed Mistral that its task was to rewrite the ground truth in its own styles.
Subsequently, we presented all questions and their corresponding ground truths from the training dataset to Mistral for style rewriting. We recorded Mistral’s predictions and paired these with the original inputs to form a new training dataset, which we named "style transferred ground truth."
We noticed that even when provided with the ground truth, the Mistral model struggles to successfully rewrite answers for mathematical tasks. Therefore, we ran the rewritten math answers through our evaluation script and found that only 73.7% of the GSM8K answers and 24% of the math algebra answers passed the script. For math tasks, we only used the answers that passed the script for training. In the ECQA and ESNLI tasks, we allowed the Mistral model to rewrite the ground truth once and directly used all the transformed ground truth answers for training.
This experiment was conducted solely with Mistral and not with the Llama model, due to our observations that Llama2-13b-chat lacks sufficient in-context following capabilities required for successful ground truth style transformation. We show a failure example in AppendixD. As a comparative experiment, we created the ’correct predicted samples only’ dataset. For each sample in this dataset, we had the Mistral model perform zero-shot generation 5 times, and using an evaluation filter, we randomly selected one correct prediction that passed the filter to be paired with the original input as the target label. If the model failed to generate a correct answer in all 5 attempts, we did not include that data in the training dataset. The construction method for the ’correct predicted samples’ and ’correct predicted samples only’ datasets is similar. The only difference is that if the sample fails to produce a correct answer, we include the ground truth in the dataset as the correct answer.
As shown in Table 4, the Mistral model trained with style-transferred ground truth data outperformed the model trained with original ground truth data and other methods in all cases except for the Math Algebra dataset, where only 24% of the training data was used. To fairly compare the training effectiveness of the ’style transferred ground truth’ and the original ground truth, we recorded the IDs of the data used by the ’style transferred ground truth’. We identified a corresponding set of 240 original ground truth data points for comparison. We found that training with only these 240 data points resulted in a 17.4% accuracy using the original ground truth, which is lower than the 19.6% achieved by the ’style transferred ground truth’. This demonstrates that data processed using the ’ground truth style transfer’ method yields better training outcomes than data labeled by humans.
Interestingly, on the ECQA and ESNLI datasets, models trained with correct answers generated by Mistral in a zero-shot fashion performed worse than those trained directly with ground truth. By reading the ’correct predicted samples only’ dataset, we found that many instances classified as correct by the evaluation filter did not result from sound reasoning by the model. Although the final answers were technically correct, the underlying rationales were incorrect. Therefore, the effectiveness of the self-training method depends significantly on whether the evaluation filter used to select correct predictions can truly identify accurate responses.
6 A hybrid approach: creating responses using “minimal change” principal
Generally speaking, our experiments highlight two essential characteristics of effective responses. First, the target LLM must be "familiar with" the response. Second, the response must be accurate. The approach detailed in Section 5.2 utilizes outputs generated by the target LLM to ensure "familiarity" and employs "rewriting ground-truth response" to maintain accuracy. Familiarity is effectively assured as the response is produced by the model itself, often resulting in low perplexity. However, as shown in the experiments in Section 5.2, the strategy of "rewriting ground-truth response" frequently fails to guarantee correctness. In this section, we explore a hybrid approach where we use responses from the target LLM to achieve high familiarity and employ GPT-4 to make minimal adjustments to correct any inaccuracies, thereby ensuring both familiarity and correctness.
We call this approach “minimal change” and its construction process is as follows: For a given question, we first request a response from the target LLM, such as Llama2 or Mistral. We then instruct GPT-4 to copy the correct portions of the prediction and revise any erroneous parts of the response. When correcting errors, GPT-4 is required to make minimal changes to accurately adjust the answer  [^1] Please note that achieving such a process requires an advanced LLM, as the LLM needs to fully understand the solution and follows the “minimal change” principal. Note that this construction process ensures that the target LLM is “familiar” with the response as the response is largely generated from itself. The “minimal change” principal, on the other hand, minimizes the influence from the GPT-4.
This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained. To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ’Minimum Change Prompt and Method’ section of the Appendix C, which includes two to three examples of minimal changes along with explanations for each modification.
Method Model Type GSM8K Math Algebra ECQA MBPP HumanEval Avg Perplexity
Groundtruth Mistral 0.434 0.206 0.68 0.276 0.323 15.03
GPT-4 Answer Directly 0.597 0.302 0.722 0.354 0.365 3.81
Minimum Change on Mistral Predictions 0.562 0.314 0.699 0.354 0.409 2.47
Minimum Change on LLaMA predictions 0.547 0.297 0.709 0.364 0.427 3.51
Groundtruth Llama 0.364 0.126 0.631 0.2 0.146 6.18
GPT-4 Answer Directly 0.428 0.150 0.656 0.2 0.158 3.58
Minimum Change on LLaMA predictions 0.433 0.166 0.649 0.2 0.213 2.75
Minimum Change on Mistral Predictions 0.402 0.161 0.647 0.218 0.183 3.32
Table 5: Comparing the experimental results of GPT4, groundtruth and minimum change
Results from our ’Minimum Change’ experiments are presented in Table 3. The data shows consistently lower perplexity compared to ground truth data across all datasets, indicating a high level of model familiarity with LLM-generated data. Furthermore, the performance across various setups was not only similar (with in-domain performance differences within 5% on most datasets) but also comparable to models trained with direct GPT-4 responses. Although the data was minimally modified based solely on the outputs from Mistral/LLaMA, theoretically, the logic and quality of the modified answers, as well as the level of detail, should be inferior to answers directly generated by GPT-4. Surprisingly, on the HumanEval and math algebra tasks, the performance of models trained with minimally changed data even surpassed those trained with data generated by GPT-4. On other datasets, the performance of the minimum change data was comparable to that of GPT-4. This underscores the effectiveness of training LLMs on data distributions familiar to them, demonstrating that minimal changes to maintain the original structure can be as effective as more detailed alterations. For more information about this experiment, please refer to the Appendix. B
7 Related Works
There are many studies related to using LLMs to generate training data for training models.
Self-Training: Several works utilize a model’s own predictions for self-training. For instance, STAR Zelikman et al. (2022) and REST Gulcehre et al. (2023) generates a dataset through sample production from the LLMs, subsequently utilizing these samples to enhance the LLMs via training. RESTem Singh et al. (2023) enhances model performance by using initial predictions, filtering for accuracy, and retraining the model with correct predictions. This iterative process improves the model’s accuracy over multiple cycles.
Data Augmentation: Numerous studies have investigated the use of synthetic data primarily for data augmentation purposes. In scenarios where sample sizes are insufficient, these studies generate additional training data to augment existing datasets (usually when the original training data is not adequate) Dai et al. (2023); Edwards et al. (2022); Møller et al. (2024); Guo et al. (2023); Ubani et al. (2023); Agrawal et al. (2023); Piedboeuf and Langlais (2023), or for generating extensive training data when high-quality human-annotated instruction tuning data is difficult to obtain Kieser et al. (2023); Taori et al. (2023); Peng et al. (2023); Xu et al. (2024). In cases where training data quality is low, many studies improve model training by providing more rationales Zhang et al. (2024); Kang et al. (2023); Li et al. (2022) or step-by-step reasoning Hsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023) to small models.
Our work distinguishes itself from existing studies by focusing on the role of familiarity in training. We explore the effectiveness of synthetic data from the perspective of familiarity, providing a theoretical foundation that substantiates the efficacy of past methods using synthetic data. Our finding offers a nuanced perspective that underscores the importance of dataset familiarity in enhancing model performance.
8 Conclusion
We systematically proved two hypotheses: LLMs are more familiar with responses generated by other LLMs when given similar inputs, and LLMs perform better on data they are familiar with. These hypotheses also explain why data annotated by LLMs yields better training outcomes than human-annotated data. The data generated by LLMs is more familiar to the trained LLMs, and training on unfamiliar data typically results in poorer outcomes. Based on this theory of familiarity, we have proposed several methods to generate data with high familiarity. Without using advanced LLMs, models can perform style transfers on ground truth to rewrite it in their own style. When using advanced LLMs for data processing, we can either use GPT-4 to read questions and generate answers in its own style, or employ our minimum change data generation method, which in some tasks, yields even better training results than direct responses from GPT-4.
9 Limitations
Perplexity can serve as a straightforward tool for measuring familiarity, but it is not very precise. It only provides relatively accurate predictions of performance decline when there are substantial differences in perplexity levels. We need a better tool to predict a model’s familiarity with a dataset and how that familiarity impacts the model’s learning.
There may be additional forms of familiarity, such as linguistic familiarity and logical familiarity. Accurately identifying and measuring the impact of these forms of familiarity presents a challenge.
Appendix A Appendix: Comparing varients of GPT-4 created training datasets
GPT-4 Answer Directly: For mathematics and coding problems, we directly present the problems from the training dataset to GPT-4 for answers. For classification tasks, we provide GPT-4 with the gold-label (which does not include human-annotated rationales) and the input, and then use its responses. We combine these answers with the original questions to create the GPT-4 Answer Training Dataset. We purposefully do not supply GPT-4 with the solutions or rationales for the math, coding, or classification problems to prevent it from merely replicating the ground truth’s problem-solving and analytical processes. If these processes were included, it could lead to a portion of GPT-4’s predictions not being generated in its own style, thus rendering our perplexity measurements—which assess how well an LLM handles predictions from other LLMs—less accurate.
Step by step transformation on GT: For classification tasks, we provide GPT-4 with both the questions and the ground truth target labels along with human-annotated rationales. For mathematical problems, we present the problems along with their answers, including the solutions process. We instruct GPT-4 to rewrite the human-annotated solutions or rationales in a step-by-step format, adding details where necessary to fill any gaps. If the existing content is already detailed enough, no additional details are added. We also instruct GPT-4 to preserve the original linguistic style as closely as possible by frequently incorporating language from the original ground truth.
Detailed Step-by-Step Transformation of Ground Truth: This data construction method is very similar to the Step by Step method. The only difference is that we specifically instruct GPT-4 to add extra details when the original ground truth lacks sufficient information.
Add details to GT: This approach involves enhancing the existing ground truth data by supplementing it with additional details to provide a richer context when necessary. We direct GPT-4 to enhance the content while maintaining the inherent language style of the original data. This method is applied selectively, only when the original information lacks sufficient clarity or depth to support effective model training and understanding.
Rewrite GT in GPT-4 style words: For this method, we present GPT-4 with the ground truth data, including human-annotated rationales and problem-solving steps. The objective is to have GPT-4 rewrite the ground truth using its own language style.
Datasets with explicit reasoning
Math Commensense Reasoning NLI
Method Training Dataset and Model Type GSM8K Math Algebra ECQA Perplexity Toekn Count
Groundtruth GSM8K, Mistral 0.434 0.162 0.594 4.72 128.882
GPT-4 Answer Directly 0.597 0.246 0.597 2.34 179.941
Rewrite GT in GPT-4 style words 0.471 0.192 0.6 4.77 156.696
Step by step 0.574 0.224 0.591 2.27 282.9
Add details to GT 0.59 0.235 0.596 3.80 272.844
Step by step transformation of GT 0.499 0.190 0.604 2.67 235.892
Detailed Step-by-Step Transformation of Ground Truth 0.506 0.196 0.602 2.59 315.028
Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.882
GPT-4 Answer Directly 0.428 0.127 0.575 2.30 179.941
Rewrite GT in GPT-4 style words 0.394 0.127 0.58 3.65 156.696
Step by step 0.402 0.134 0.553 2.19 290.2
Add details to GT 0.38 0.128 0.579
Step by step transformation of GT 0.394 0.146 0.573 2.42 235.892
Detailed Step-by-Step Transformation of Ground Truth 0.396 0.114 0.553 2.37 315.028
Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.93 163.355
GPT-4 Answer Directly 0.553 0.301 0.608 2.20 328.764
Step by step 0.533 0.239 0.583 2.25 452
Add details to GT 0.445 0.230 0.57 3.148 417.35
Step by step transformation of GT 0.444 0.210 0.577 2.65 370.349
Detailed Step-by-Step Transformation of Ground Truth 0.468 0.223 0.582 2.62 484.368
Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.17 163.355
GPT-4 Answer Directly 0.35 0.150 0.561 2.591 328.764
Step by step 0.344 0.144 0.593 2.55 450.1
Add details to GT 0.38 0.128 0.579 3.31 417.35
Step by step transformation of GT 0.376 0.140 0.565 2.80 370.349
Detailed Step-by-Step Transformation of Ground Truth 0.299 0.110 0.545 2.816 484.368
Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.28 73.508
GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.536
Rewrite GT in GPT-4 style words 0.384 0.194 0.721 9.26 134.338
Step by step 0.481 0.203 0.71 4.66 325.2
Add details to GT 0.422 0.204 0.731 9.08 238.292
Step by step transformation of GT 0.462 0.194 0.681 6.858 218.74
Detailed Step-by-Step Transformation of Ground Truth 0.487 0.186 0.68 5.61 322.855
Groundtruth ECQA, Llama 0.132 0.079 0.631 12.16 73.508
GPT-4 Answer Directly 0.379 0.155 0.656 5.33 176.536
Rewrite GT in GPT-4 style words 0.282 0.115 0.664 6.84 134.338
Step by step 0.363 0.140 0.648 4.46 337.3
Add details to GT 0.062 0.097 0.663 8.47 238.292
Step by step transformation of GT 0.363 0.139 0.648 5.54 218.748
Detailed Step-by-Step Transformation of Ground Truth 0.135 0.106 0.66 5.08 322.855
Table 6: Appendix A: Additional experiments for Table2
Appendix B Appendix: Minimum Change Vs GPT4 answer
Datasets with explicit reasoning
Math Commensense Reasoning NLI Code
Method Training Dataset and Model Type GSM8K Math Algebra ECQA ESNLI HumanEval MBPP Perplexity Toekn Count
Groundtruth GSM8K, Mistral 0.434 0.162 0.594 0.597 4.53 128.88
GPT-4 Answer Directly 0.597 0.246 0.597 0.587 2.34 179.941
Mistral Minimum Change Data 0.562 0.234 0.597 0.558 1.67 177.156
llama minimum change data 0.547 0.242 0.621 3.19 166.152
Groundtruth GSM8K, Llama 0.371 0.143 0.566 0.559 3.27 130.856
GPT-4 Answer Directly 0.428 0.127 0.575 0.568 2.30 183.852
Llama Minimum Change Data 0.433 0.140 0.602 0.602 2.22 166.15
Mistral Minimum Change Data 0.402 0.141 0.576 2.249 177.156
Groundtruth Math algebra, Mistral 0.264 0.206 0.554 0.574 5.83 163.355
GPT-4 Answer Directly 0.553 0.301 0.608 0.555 2.19 328.764
Mistral Minimum Change Data 0.536 0.313 0.622 0.558 1.85 297.609
llama minimum change data 0.546 0.296 0.63 2.47 269.688
Groundtruth Math algebra, Llama 0.36 0.126 0.575 0.552 5.14 162.417
GPT-4 Answer Directly 0.35 0.150 0.561 0.564 2.59 326.573
Llama Minimum Change Data 0.412 0.166 0.602 0.554 2.268 269.68
Mistral Minimum Change Data 0.32 0.160 0.568 2.53 297.609
Groundtruth ECQA, Mistral 0.258 0.1348 0.68 0.544 51.29 73.508
GPT-4 Answer Directly 0.462 0.223 0.722 0.589 6.35 176.53
Mistral Minimum Change Data 0.433 0.192 0.699 0.542 3.798 112.591
LLama Minimum Change Data 0.483 0.212 0.709 4.89 128.695
Groundtruth ECQA, Llama 0.132 0.079 0.631 0.431 11.95 76.118
GPT-4 Answer Directly 0.379 0.155 0.656 0.479 5.33 183.182
Llama Minimum Change Data 0.392 0.128 0.649 0.522 3.619 128.695
Mistral Minimum Change Data 0.28 0.126 0.647 5.58 112.591
Groundtruth ESNLI, Mistral 0.417 0.172 0.594 0.814 378.7 29.925
GPT-4 Answer Directly 0.464 0.214 0.592 0.804 6.54 137.967
Mistral Minimum Change Data 0.453 0.206 0.602 0.831 3.38 82.367
Llama Minimum Change Data 0.432 0.203 0.63 0.836 4.15 97.078
Groundtruth ESNLI, Llama 0.332 0.125 0.582 0.751 1201.70 29.963
GPT-4 Answer Directly 0.386 0.123 0.564 0.776 5.28 143.739
Llama Minimum Change Data 0.357 0.128 0.578 0.796 3.74 97.078
Mistral Minimum Change Data 0.346 0.130 0.546 0.811 4.34 82.367
Groundtruth HumanEval, Mistral 0.362 0.191 0.583 0.531 0.323 5.69 71.60
GPT-4 Answer Directly 0.313 0.162 0.581 0.442 0.365 4.71 107.9
Mistral Minimum Change Data 0.4055 0.202 0.5915 0.541 0.408 2.21 109.073
Llama Minimum Change Data 0.4115 0.179 0.5835 0.5505 0.426 2.451 122.0
Groundtruth HumanEval, Llama 0.0705 0.083 0.528 0.339 0.146 5.75 70.8
GPT-4 Answer Directly 0.1245 0.105 0.553 0.473 0.158 4.33 107.14
Llama Minimum Change Data 0.2445 0.389 0.5705 0.5175 0.213 2.21 122.0
Mistral Minimum Change Data 0.131 0.103 0.528 0.4235 0.183 2.75 109.07
Groundtruth MBPP, Mistral 0.392 0.175 0.527 0.481 0.276 7.78 71.60
GPT-4 Answer Directly 0.399 0.186 0.568 0.553 0.354 3.419 786
Mistral Minimum Change Data 0.405 0.179 0.601 0.48 0.354 2.817 63.86
Llama Minimum Change Data 0.332 0.179 0.56 0.511 0.364 4.55 54.96
Groundtruth MBPP, Llama 0.328 0.131 0.556 0.537 0.2 4.74 73.90
GPT-4 Answer Directly 0.351 0.139 0.574 0.553 0.202 3.33 69.78
Llama Minimum Change Data 0.34 0.130 0.569 0.539 0.204 3.44 54.96
Mistral Minimum Change Data 0.333 0.123 0.563 0.544 0.218 3.45 63.86
Table 7: Appendix B: More results for Table 5. Comparing the experimental results of GPT4, groundtruth and minimum change
Appendix C Appendix: Minimum Change Prompt and Method
Figure 3: Appendix C: Minimum Change Data Correction Examples
Its construction process is as follows: For a given question, we first ask the target LLM, such as Llama2 or Mistral, to produce a response. If the response is correct, we include it in the synthetic dataset. If not, we instruct GPT-4 to revise it, making minimal changes to the original answer to correct it  [^2] Please note that achieving such a process requires an advanced LLM, as the LLM needs to fully understand the solution and follows the “minimal change” principal. Note that this construction process ensures that the target LLM is “familiar” with the response as the response is largely generated from itself. The “minimal change” principal, on the other hand, minimizes the influence from the GPT-4.
Specifically, minimal change is implemented via the following steps:
  • Initial Prediction the model generates an initial prediction.
  • Minimal Modification uses GPT-4 to adjusts the initial prediction with the fewest possible changes to correct errors, deliberately preserving the original distribution by keeping as many words unchanged as possible. An illustration of such a modification is provided in Figure 2. We have detailed the prompt used to guide GPT-4 in the ’Minimum Change Prompt’ section of the Appendix.
  • Training with Modified Predictions the minimum changed predictions are subsequently used as target labels for model training.
Figure 4: Appendix C: Minimum Change Prompt Example
This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained. To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ’Minimum Change Prompt’ section of the Appendix, which includes two to three examples of minimal changes along with explanations for each modification.
Appendix D Appendix: LLaMA groundtruth style transfer failure example
The LLaMA-13b-chat model often fails to follow our in-context examples, leading to incorrect outputs when it is tasked with modifying ground truth data.
Figure 5: Appendix D: LLaMA groundtruth style transfer failure example

### v3

@@
 Abstract
-This paper explores an intriguing observation: fine-tuning a large language model (LLM) with responses generated by a LLM often yields better results than using responses generated by humans.
+This paper explores an intriguing observation: fine-tuning a large language model (LLM) with responses generated by a LLM often yields better results than using responses generated by humans, particularly in reasoning tasks.
 We conduct an in-depth investigation to understand why this occurs.
-Contrary to the common belief that these instances is simply due to the more detailed nature of LLM-generated content, our study identifies another contributing factor: an LLM is inherently more "familiar" with LLM generated responses.
+Contrary to the common belief that these instances is due to the more detailed nature of LLM-generated content, our study identifies another contributing factor: an LLM is inherently more "familiar" with LLM generated responses.
 This familiarity is evidenced by lower perplexity before fine-tuning.
 We design a series of experiments to understand the impact of the "familiarity" and our conclusion reveals that this "familiarity" significantly impacts learning performance.
-Training with LLM-generated responses not only enhances performance but also helps maintain the model’s capabilities in other tasks after fine-tuning on a specific task.
+Training with LLM-generated responses not only enhances performance but also helps maintain the model’s capabilities in other reasoning tasks after fine-tuning on a specific task.
+Our code and data are public at https://github.com/XuanRen4470/I-Learn-Better-If-You-Speak-My-Language.
 I Learn Better If You Speak My Language: Understanding the Superior Performance of Fine-Tuning Large Language Models with LLM-Generated Responses
-Xuan Ren xuan.ren@adelaide.edu.au                        Biao Wu biaowu165534@gmail.com                        Lingqiao Liuthanks: Corresponding author. lingqiao.liu@adelaide.edu.au
+Xuan Ren University of Adelaide xuan.ren@adelaide.edu.au                        Biao Wu University of Technology Sydney biaowu165534@gmail.com                        Lingqiao Liuthanks:   Corresponding author.
+University of Adelaide lingqiao.liu@adelaide.edu.au
 1 Introduction
@@
 (2023); Ubani et al.
-(2023); Agrawal et al.
+(2023); Piedboeuf and Langlais (2023a); Agrawal et al.
 (2023); Xu et al.
@@
 (2023).
-The quality of data generated by large language models (LLMs) is considered to be very high and conducive to effective trainingLee et al.
-(2023).
-The prevailing view often attributes the successful training results of LLM-generated data to the fact that LLMs provide more details than human-annotated data, such as ’chain-of-thought-style’ reasoning processHsieh et al.
+In this paper, we made an intriguing observation that fine-tuning an LLM (dubbed target LLM) with responses generated by a LLM, either target LLM or other LLMs, often yields better results than using responses provided by humans in chain-of-thought reasoning tasks, as shown in Table 1.
+The prevailing view attributes the successful training results of LLM-generated data to the fact that LLMs provide more details than human-annotated data, such as ‘chain-of-thought-style’ reasoning processHsieh et al.
 (2023); Ho et al.
@@
 However, our experiments have found numerous counterexamples.
-As shown in Figure 1, target labels that include more details and utilize chain of thought styles do not necessarily result in better training outcomes.
-Through experiments on various tasks, we have observed that when LLMs directly generate target labels, the training results are often superior to those obtained by transforming ground truth into more detailed, step-by-step training.
-This indicates that whether the answer generation is executed in the LLM’s own style or is based on a rewritten human style is a key factor affecting training outcomes.
-Figure 1: This figure illustrates the training outcomes associated with various data generation methods.
-As demonstrated by the examples shown, the inclusion of additional details does not necessarily translate into enhanced performance.
-This observation motivates us to investigate the underlying reason for this phenomenon. We found that whether the LLM is familiar with the target label is a crucial factor influencing model training.
+As shown in Figure 1 , target responses that include more details and utilize chain of thought styles do not necessarily result in better training outcomes.
+This indicates there are other factors contributing to the excellent performance of LLM generated responses.
+Figure 1: This figure shows training outcomes for different data generation methods, demonstrating that more details do not always improve performance.
+The observation above motivates us to investigate the underlying reason for this phenomenon.
+We found that whether the target LLM is familiar with the target responses is a crucial factor influencing model training on reasoning tasks.
 Our main hypothesis are as follows:
-  • Given the same question, an LLM is inherently familiar with responses generated by itself or other LLMs.
-  • an LLM can better adapt to and leverage responses in its familiar styles, leading to improved outcomes.
+  • Given the same question, the target LLM is inherently familiar with responses generated by itself or other LLMs.
+  • The target LLM can better adapt to and leverage responses in its familiar styles, leading to improved outcomes.
 The following section of this paper will examine the hypothesis through a series of experiments.
-We will measure ’familiarity,’ conduct ablation studies by creating different variants of LLM-generated responses, and evaluate their performance to understand various factors in the process.
+We will measure ‘familiarity,’ conduct ablation studies by creating different variants of LLM-generated responses, and evaluate their performance to understand various factors in the process.
 This will provide insights into our hypothesis.
 2 Empirical Study Protocols
-The purpose of this paper is to understand and investigate the aforementioned phenomenon rather than to propose a new method.
-Our study’s methodology is based on empirical analysis.
+The purpose of this paper is to understand and investigate the aforementioned phenomenon through empirical analysis.
 We begin by introducing our study protocols, including the datasets and evaluation approaches used.
-2.1 Evaluation approaches
+2.1 Training and Evaluation approaches
 Without loss of generality, we denote the training data as question and response pairs $(q,r)$.
@@
 (2023) to create the training dataset.
-We will use a subset of the LLaMA2 benchmark [34] that contains human-labeled explicit reasoning processes to create the training dataset.
+We will use a subset of the Llama2 benchmark that contains human-labeled explicit reasoning processes to create the training dataset.
 The question and response pairs in these datasets are human-constructed.
@@
 (2023) models involves specific adaptations to optimize performance.
-Notably, the warm-up period is set to 10% of the total training steps consistently across our experiments.
-For the LLaMA-13b-chat model training on the HumanEval dataset, we set the learning rate to 2e-4.
+For general, fully supervised fine-tuning of the Llama2-13b-chat and Mistral-7b-Instruct-v2 models, we employ a default learning rate of 2e-5, a batch size of 32, and a warm-up period comprising 10% of the total steps.
+In contrast, Given the small size of the training datasets, such as MBPP and HumanEval, we reduced the batch size to 10 to accommodate the limited data.
+For the Llama2-13b-chat model training on the HumanEval dataset, we set the learning rate to 2e-4.
 This adjustment was necessary as lower rates proved insufficient for the model to effectively learn from this dataset.
-Given the small size of the training datasets, such as MBPP and HumanEval, we reduced the batch size to 10 to accommodate the limited data.
-In contrast, for general fine-tuning of the LLaMA2-13b-chat and Mistral-7b-Instruct-v2 models, we employ a default learning rate of 2e-5, a batch size of 32, and a warm-up period comprising 10% of the total steps.
 Additionally, we use a cosine learning rate schedule and fine-tune only the Q and V matrices of the LORAHu et al.
@@
 The other is called cross-task evaluation, which evaluates performance across tasks on other datasets.
-The purpose of using cross-domain evaluation is to test if the model can benefit from the training process of a similar task and if the model will forget its capability on other tasks after training on a specific task.
-3 Experiment setup
-3.1 Dataset and Evaluation
+The purpose of using cross-task evaluation is to test if the model can benefit from the training process of a similar task and if the model will forget its capability on other tasks after training on a specific task.
+We have developed an evaluation script to automatically measure and report the pass@1 accuracy of models for each dataset.
+For the HumanEval dataset, we utilize its built-in testing script to ensure the accuracy of results.
+Our evaluation scripts require answers to follow a specific format, detailed in Appendix F.
+The models not trained in this format have demonstrated difficulty adhering to it for coding tasks, such as MBPP and HumanEval Chen et al.
+(2021); Austin et al.
+(2021).
+Therefore, cross-domain evaluations are not conducted for MBPP and HumanEval.
+2.2 Dataset and API
 We utilized the Llama2 benchmark Touvron et al.
-(2023) to select five tasks that contain human-labeled explicit reasoning.
-These were evaluated based on accuracy to assess the effectiveness of different data construction methods.
-The term ’dataset with explicit reasoning’ refers to datasets that include a detailed reasoning process in their annotations.
+(2023) to select five datasets that include a detailed reasoning process in their annotations.
 This category includes GSM8KCobbe et al.
@@
 Our selection includes all coding and mathematics datasets within the Llama2 benchmark.
-We chose to use the ECQA dataset because, unlike the CQATalmor et al.
-(2018) dataset included in the LLaMA2 benchmark, ECQA includes rationales provided by human annotators.
+We chose to use the ECQA dataset because it includes human annotated rationales, unlike the CQATalmor et al.
+(2018) dataset included in the Llama2 benchmark.
 Code: We evaluated coding ability using the HumanEval Chen et al.
@@
 (2021) datasets.
-We report the pass@1 accuracy for each dataset.
 Note that the HumanEval dataset contains only 164 testing data points with no training data.
@@
 As a result, we obtained 1000 training and 1314 testing data points for GSM8K, and 1000 training and 752 testing data points for MATH (algebra).
-Evaluation Script:
-To facilitate automatic evaluation across all models and datasets, we developed a specialized prompt format requiring models to conclude their predictions with ’Final Answer: answer.’ For instance, the output should read ’Final Answer: 100’ for an answer of 100.
-Our evaluation script extracts the text following the ’Final Answer:’ substring.
-For the ECQA task, it verifies if the first text after ’Final Answer:’ correctly matches one of the answer options.
-For mathematical problems including the GSM8K and MATH datasets, the script checks if the number after the substring ’Final Answer:’ aligns with the ground truth number.
-We utilize the built-in testing script for the HumanEval dataset to ensure result accuracy and employ our custom scripts for all other tasks.
-By adapting the training data’s question and answer formats to this style, our models maintain consistent response formatting across different tasks.
-The evaluation scripts for HumanEval Chen et al.
-(2021) and MBPP Austin et al.
-(2021) require answers in a specific format.
-Models that have not been trained on this format struggle to follow it correctly, even when provided with in-context examples.
-Consequently, we do not conduct cross-domain evaluations on the MBPP and HumanEval datasets.
-4 Observations and Hypothesis:
-4.1 Observation 1: Superior Performance of GPT-4 Over Human-Annotated Labels
-Method Training Dataset and Model Type GSM8K Math Algebra ECQA HumanEval MBPP
-Groundtruth GSM8K, Mistral 0.434 0.162 0.594
-GPT-4 Answer Directly 0.597 0.246 0.597
-Groundtruth GSM8K, Llama 0.364 0.141 0.565
-GPT-4 Answer Directly 0.428 0.128 0.575
-Groundtruth Math algebra, Mistral 0.264 0.206 0.554
-GPT-4 Answer Directly 0.553 0.302 0.608
-Groundtruth Math algebra, Llama 0.36 0.126 0.575
-GPT-4 Answer Directly 0.35 0.150 0.561
-Groundtruth ECQA, Mistral 0.258 0.134 0.68
-GPT-4 Answer Directly 0.462 0.223 0.722
-Groundtruth ECQA, Llama 0.132 0.0798 0.631
-GPT-4 Answer Directly 0.379 0.156 0.656
-Groundtruth HumanEval, Mistral 0.3625 0.191 0.583 0.323
-GPT-4 Answer Directly 0.313 0.163 0.581 0.366
-Groundtruth HumanEval, Llama 0.0705 0.083 0.528 0.146
-GPT-4 Answer Directly 0.1245 0.105 0.552 0.159
-Groundtruth MBPP, Mistral 0.392 0.176 0.527 0.276
-GPT-4 Answer Directly 0.399 0.186 0.568 0.354
-Groundtruth MBPP, Llama 0.328 0.132 0.556 0.2
-GPT-4 Answer Directly 0.351 0.140 0.574 0.202
-Table 1: Comparison of human-annotated data versus data generated directly by GPT-4.
-We employ GPT-4 to generate responses using only the questions from the training datasets.
-For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation.
-In-domain performance is highlighted in grey.
-Data points significantly lower in accuracy by 15% or more relative to another method are marked in red.
-Our study is inspired by a compelling observation: when a sophisticated LLM, such as GPT-4, is used to generate responses to questions, and these questions along with their generated responses are used as a synthetic dataset, fine-tuning smaller LLMs (we use both Llama2-13B-Chat and Mistral-7B-Instruct-v2) on this synthetic dataset often yields better performance than using the original dataset with human-provided answers.
+API: We did our experiment with use gpt-4-1106-preview OpenAI (2023) from OpenAI and claude-3-5-sonnet-20240620 Anthropic (2023) from Anthropic.
+3 Observation and Hypothesis
+3.1 Observation 1: Superior Performance of LLMs Over Human-Annotated Labels
+Our study is inspired by a compelling observation: when an advanced LLM, such as GPT-4 or Claude 3.5, is used to generate responses to questions, and these questions along with their generated responses are used as a synthetic dataset, fine-tuning a smaller LLM (such as Llama2-13B-Chat or Mistral-7B-Instruct-v2) on this synthetic dataset often yields better performance than using the original dataset with human-provided answers.
 The detailed experimental results are presented in Table 1.
-As demonstrated, the majority of the low-performance data points (14 out of 15) marked in red reflect accuracies when training with human-annotated ground truth.
+As demonstrated, the majority of the low-performance data points (14 out of 17), marked in red, reflect accuracies when training with human-annotated ground truth.
 Using synthetic datasets consistently yields higher in-domain performance compared to human-answered datasets.
@@
 This observation motivates us to investigate the underlying reasons.
-4.2 Observation 2: significantly lower perplexity of GPT-4 and other LLMs Over Human-Annotated Labels
+3.2 Observation 2: Significantly Lower Perplexity of LLM generated responses Over Human-Annotated Labels
 Figure 2: Average Perplexity Comparison
-We measure the perplexity of responses as assessed by the target (smaller) LLMs.
-An interesting observation, shown in Figure 2, is that, given the input questions from the responses generated by GPT-4 attain significantly lower perplexity compared to those provided by humans.
-To explore whether this phenomenon is unique to GPT-4 or common among LLM responses, we employed a zero-shot approach with the Mistral-7B-Instruct-v2 model and the LLaMA-13B-chat model.
-These models answered the same questions from the training set.
-As shown in Figure 2, their perplexity was also significantly lower than that of the human-labeled dataset.
-Lower perplexity indicates that LLMs are more familiar with responses generated by themselves or other LLMs than with human-generated responses.
+We measured the perplexity of responses as assessed by the target LLMs (Mistral-7B-Instruct-v2 and Llama2-13B-chat).
+An interesting observation, shown in Figure 2, is that responses generated by LLMs such as GPT-4, Claude 3.5, Mistral-7B-Instruct-v2, and Llama2-13B-chat, using a zero-shot approach, exhibit significantly lower perplexity compared to those provided by humans for the same input questions from the training data.
+Lower perplexity suggests that LLMs are more familiar with responses generated by other LLMs or themselves than with human-generated responses.
 3.3 Hypotheses
 From Observation 2, we notice that given the same question, the average perplexity of responses generated by LLMs is consistently lower than the perplexity of ground truth annotated by humans.
-This observation leads us to pose our first hypothesis: When generating target labels for the same question, one LLM inherently more familiar with data generated by another LLM.
-We not only observed consistently lower perplexity from GPT-4 across tasks in Observation 2 but also noted that LLMs trained on GPT-4 generated responses consistently outperform the same LLMs trained using human-annotated data.
+This observation leads us to pose our first hypothesis: When generating target responses for the same question, one LLM inherently more familiar with data generated by another LLM.
+We not only observed consistently lower perplexity from advanced LLMs (GPT-4 and Claude 3.5) across tasks in Observation 2 but also noted that LLMs trained on advanced LLMs generated responses consistently outperform the same LLMs trained using human-annotated data.
 This phenomenon leads us to propose our second hypothesis: An LLM performs better when trained on data it is familiar with.
-5 Investigation of the proposed hypotheses
-Method Training Dataset and Model Type GSM8K Math Algebra ECQA Perplexity token length
-Groundtruth GSM8K, Mistral 0.434 0.162 0.594 4.54 128.88
-GPT-4 Answer Directly 0.597 0.246 0.597 2.35 179.94
-Step by step 0.574 0.224 0.591 2.27 282.9
-Step by step transformation of GT 0.499 0.190 0.6 2.67 235.892
-Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.88
-GPT-4 Answer Directly 0.428 0.128 0.575 2.30 179.941
-Step by step 0.402 0.134 0.553 2.19 290.92
-Step by step transformation of GT 0.394 0.146 0.573 2.43 242.22
-Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.83 163.36
-GPT-4 Answer Directly 0.553 0.302 0.608 2.20 328.76
-Step by step 0.533 0.239 0.583 2.249 451.99
-Step by step transformation of GT 0.444 0.210 0.577 2.65 370.35
-Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.15 163.36
-GPT-4 Answer Directly 0.35 0.150 0.561 2.59 328.76
-Step by step 0.344 0.145 0.593 2.55 450.08
-Step by step transformation of GT 0.376 0.141 0.565 2.81 369.06
-Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.3 73.51
-GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.54
-Step by step 0.481 0.203 0.71 4.66 325.15
-Step by step transformation of GT 0.462 0.194 0.681 6.89 218.748
-Groundtruth ECQA, Llama 0.132 0.0798 0.631 11.96 73.51
-GPT-4 Answer Directly 0.379 0.156 0.656 5.33 176.54
-Step by step 0.363 0.140 0.648 4.46 337.30
-Step by step transformation of GT 0.312 0.108 0.628 5.55 226.99
-Table 2: Performance comparison of models trained on data constructed by different methods across various tasks.
-Data points are labeled as low performance when their accuracy is more than 15% lower relative to the highest accuracy achieved in the same dataset using the same model.
+Method Training Dataset and Model Type GSM8K Math Algebra ECQA HumanEval MBPP
+Groundtruth GSM8K, Mistral 0.434 0.162 0.594
+GPT-4 Answer Directly 0.597 0.246 0.597
+Claude Answer Directly 0.586 0.230 0.595
+Groundtruth GSM8K, Llama 0.364 0.141 0.565
+GPT-4 Answer Directly 0.428 0.128 0.575
+Claude Answer Directly 0.433 0.110 0.548
+Groundtruth Math algebra, Mistral 0.264 0.206 0.554
+GPT-4 Answer Directly 0.553 0.302 0.608
+Claude Answer Directly 0.554 0.277 0.606
+Groundtruth Math algebra, Llama 0.36 0.126 0.575
+GPT-4 Answer Directly 0.35 0.150 0.561
+Claude Answer Directly 0.317 0.137 0.54
+Groundtruth ECQA, Mistral 0.258 0.134 0.68
+GPT-4 Answer Directly 0.462 0.223 0.722
+Claude Answer Directly 0.457 0.213 0.714
+Groundtruth ECQA, Llama 0.132 0.0798 0.631
+GPT-4 Answer Directly 0.379 0.156 0.656
+Claude Answer Directly 0.38 0.129 0.678
+Groundtruth HumanEval, Mistral 0.363 0.191 0.583 0.323
+GPT-4 Answer Directly 0.313 0.163 0.581 0.366
+Claude Answer Directly 0.383 0.181 0.553 0.372
+Groundtruth HumanEval, Llama 0.0705 0.083 0.528 0.146
+GPT-4 Answer Directly 0.125 0.105 0.552 0.159
+Claude Answer Directly 0.296 0.104 0.435 0.140
+Groundtruth MBPP, Mistral 0.392 0.176 0.527 0.276
+GPT-4 Answer Directly 0.399 0.186 0.568 0.354
+Claude Answer Directly 0.439 0.189 0.573 0.37
+Groundtruth MBPP, Llama 0.328 0.132 0.556 0.2
+GPT-4 Answer Directly 0.351 0.140 0.574 0.202
+Claude Answer Directly 0.339 0.122 0.546 0.204
+Table 1: Human-annotated data Vs.
+data generated directly by GPT-4/Claud 3.5.
+In-domain performance is highlighted in grey.
+Data points are highlighted when accuracy is more than 15% below the highest accuracy on the same dataset and model.
+There are 14, 1, and 2 red data points for Groundtruth, GPT-4, and Claude, respectively.
+4 Investigation of the hypotheses
+Method Training Dataset and Model Type GSM8K Math Algebra ECQA token length
+Zeroshot Mistral 0.413 0.185 0.504
+Zeroshot Llama2 0.35 0.125 0.597
+GPT-4 Answer Directly Mistral 0.597 0.246 0.597 179.94
+GPT-4 Step-by-step 0.574 0.224 0.591 282.9
+GPT-4 Detailed Step-by-Step transformation of GT 0.506 0.196 0.602 315.028
+GPT-4 Answer Directly Llama 0.428 0.128 0.575 179.941
+GPT-4 Step-by-step 0.402 0.134 0.553 290.92
+GPT-4 Detailed Step-by-Step transformation of GT 0.396 0.114 0.553 315.028
+GPT-4 Answer Directly Mistral 0.553 0.302 0.608 328.76
+GPT-4 Step-by-step 0.533 0.239 0.583 451.99
+GPT-4 Detailed Step-by-Step transformation of GT 0.468 0.223 0.582 484.368
+GPT-4 Answer Directly Llama 0.35 0.150 0.561 328.76
+GPT-4 Step-by-step 0.344 0.145 0.593 450.08
+GPT-4 Detailed Step-by-Step transformation of GT 0.299 0.110 0.545 484.368
+GPT-4 Answer Directly Mistral 0.462 0.223 0.722 176.54
+GPT-4 Step-by-step 0.481 0.203 0.71 325.15
+GPT-4 Detailed Step-by-Step transformation of GT 0.487 0.186 0.68 322.855
+GPT-4 Answer Directly Llama 0.379 0.156 0.656 176.54
+GPT-4 Step-by-step 0.363 0.140 0.648 337.30
+GPT-4 Detailed Step-by-Step transformation of GT 0.135 0.106 0.66 322.855
+Table 2: Performance comparison of models trained on data constructed using different methods.
+$n_{\text{train}}=1000$.
+Data points are labeled as low performance when accuracy is more than 15% below the highest accuracy on the same dataset using the same model.
+See Table 6 and Table 7 for additional experiments with GPT-4 and Claude 3.5.
 In this section, we conduct an in-depth investigation about the aforementioned hypotheses.
@@
 Experiment 1: Is observation 1 only rooted from more detailed reasoning steps provided by the LLM other than familiarity?
-Our observation 1 suggests that using GPT-4 generated data leads to better training effectiveness than using human-annotated ground truth.
+Our observation 1 suggests that using advanced LLMs generated data leads to better training effectiveness than using human-annotated ground truth.
 Our second hypothesis posits that this enhanced learning by LLMs on LLMs responses could be attributed to familiarity with the data.
-It is a common belief that GPT-4 often includes more detailed content than human annotators, which could contribute to improved training outcomes.
-However, we question whether the superior results from GPT-4 are solely due to the addition of more details, or if familiarity plays a significant role as well.
-To explore whether familiarity is a crucial factor, we designed the following experiment.
+It is a common belief that LLMs often includes more detailed content than human annotators, which could contribute to improved training outcomes.
+However, we question whether the superior results from advanced LLMs are solely due to the addition of more details, or if familiarity plays a significant role as well.
+To investigate whether the superior performance is due to the additional detail provided by advanced LLMs compared to humans, we designed the following experiment.
 We employed different methods to have GPT-4 generate data in a detailed, step-by-step manner.
@@
 In addition, it is noteworthy that using the Mistral 7B model for training on math algebra, converting ground truth into step-by-step answers was even detrimental.
-Both of the step by step groundtruth transformation methods achieves accuracy below 24% on Math Algebra, much lower than 30.2% achieved by ’GPT-4 answer directly’(Gpt-4 answer directly represents employing GPT-4 to generate responses using only the questions from the training datasets.
-For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation.) We speculate that one possible reason for this is that adding ’more than enough details’ may sometimes complicate the model’s reasoning process.
-We have included additional experiments involving various GPT-4 generated variants in the appendixA, which generally conform to the patterns described above.
+Both of the step by step groundtruth transformation methods achieves accuracy below 24% on Math Algebra, much lower than 30.2% achieved by ‘GPT-4 answer directly’(Gpt-4 answer directly represents employing GPT-4 to generate responses using only the questions from the training datasets.
+For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation.) We speculate that one possible reason for this is that adding ‘more than enough details’ may sometimes complicate the model’s reasoning process.
+We have included additional experiments in the appendix A, involving various GPT-4 (refer to Table 6) and Claude 3.5 (refer to Table 7) generated variants with different token lengths (shorter or longer), which generally conform to the patterns described above.
 For clarity and conciseness, Table 2 only includes 2 examples of the step-by-step variants.
-After preliminary validation of our hypotheses, we made another intriguing observation: According to Table 2, the ’GPT-4 answer directly’ dataset, characterized by a shorter average token length of target labels compared to both ’step by step’ and ’step by step transformation of GT’ datasets, suggests fewer details in its content.
+After preliminary validation of our hypotheses, we made another intriguing observation: According to Table 2, the ‘GPT-4 answer directly’ dataset, characterized by a shorter average token length of target responses compared to both ‘step by step’ and ‘step by step transformation of GT’ datasets, suggests fewer details in its content.
 Despite this, the performance of the data directly generated by GPT-4 frequently ranks among the best of all GPT-4 answer variants.
 Meanwhile, other data generation methods exhibited significantly lower performance than GPT-4’s direct answers on at least two tasks, as highlighted in red in Table 2.
-This suggests that, given a problem, the data directly generated by GPT-4 might yield the good training results.
-Artificially defined prompts that add information to the ground truth may lead to the generation of target labels that do not conform to the LLM’s linguistic or logical preferences.
-Experiment 2: Direct measuring the impact of familiarity
+This suggests that, given a problem, the data directly generated by an advanced LLM might yield the good training results.
+We observed the same pattern in Table 6 and Table 7.
+Artificially defined prompts that add information to the ground truth may lead to the generation of target responses that do not conform to the LLM’s linguistic or logical preferences.
+Experiment 2: The impact of familiarity
 Method Model Type GSM8K GSM8K Perplexity Math Algebra Math Algebra Perpleity ECQA ECQA Perplexity
-GPT-4 answer with lower perplexity Mistral 0.597 2.35 0.301 2.20 0.722 6.35
+GPT-4 Answer with Lower Perplexity Mistral 0.600 2.20 0.303 2.10 0.722 6.15
 GPT-4 Answer with Higher Perplexity 0.547 5.58 0.231 4.66 0.702 11.13
-GPT-4 answer with lower perplexity LLaMA 0.428 2.30 0.150 2.59 0.656 5.33
+GPT-4 Answer with Lower Perplexity Llama2 0.424 2.21 0.153 2.36 0.655 5.11
 GPT-4 Answer with Higher Perplexity 0.38 4.19 0.131 4.35 0.654 7.69
-Table 3: Comparison between the GPT-4 answer with higher perplexity Vs. lower perplexity
-This section introduces a method for assessing the influence of style ’familiarity’.
-We direct GPT-4 to produce several varient of responses to the same query, grouping them into two distinct datasets based on their perplexity levels—high and low.
+Claude Answer with Lower Perplexity Mistral 0.586 2.36 0.277 2.24 0.714 4.02
+Claude Answer with Higher Perplexity 0.494 5.96 0.218 8.98 0.7 10.63
+Claude Answer with Lower Perplexity Llama2 0.433 2.04 0.137 2.33 0.678 3.51
+Claude Answer with Higher Perplexity 0.326 3.36 0.121 5.51 0.619 5.69
+Table 3: GPT-4/Claude 3.5: Answers with Lower Perplexity vs. Higher Perplexity. $n_{\text{train}}=1000$.
+This section introduces a method for assessing the influence of style ‘familiarity’.
+We direct advanced LLMs to produce several varient of responses to the same query, grouping them into two distinct datasets based on their perplexity levels—high and low.
 We then evaluate and compare the learning outcomes using these two sets.
-One concern of this experiment is that GPT-4 may produce answers with different level of details.
-To eliminate this factor, we first let GPT-4 to generate one answer and then explicitly direct it to paraphrase the generated answer.
+One concern of this experiment is that advanced LLMs may produce answers with different level of details.
+To eliminate this factor, we first let advanced LLMs to generate one answer and then direct it to paraphrase the generated answer.
 In this way, we can create responses that vary in perplexity but maintain the same semantic meaning.
@@
 As seen, training on the higher-perplexity training set consistently shows slightly lower performance than the lower perplexity counterpart, particularly noticeable in the in-domain performance on two math datasets.
-For example, in the algebra dataset from Mistral, the lower perplexity set achieves a 30% accuracy rate, while the higher perplexity set only reached 23%.
-Since they are only differed in the style of the text, this result demonstrated that the familiarity of LLMs with the stylistic aspects of target labels significantly affects training outcomes.
+For example, when training the Mistral model on the Math Algebra dataset, the GPT-4 answer with lower perplexity set achieves a 30% accuracy rate, while the GPT-4 answer higher perplexity set only reached 23%.
+Since they are only differed in the style of the text, this result demonstrated that the familiarity of LLMs with the stylistic aspects of target responses significantly affects training outcomes.
 From these findings, we conclude that LLMs perform poorly when trained on unfamiliar data, explaining the suboptimal training results with human-annotated data.
 Discussion: Why are LLMs often familiar with responses generated from other LLMs?
-The familiarity observed when training on data generated by GPT-4 might stems from the overlap in training data used during the pre-training stages of these models.
-Many LLMs, including GPT-4, share similar pre-training data, developing an understanding of specific output styles and content.
+Huh et al.
+(2024) propose the Platonic Representation Hypothesis, suggesting that different AI models exhibit convergent behavior.
+This hypothesis supports our findings that LLMs possess intrinsic familiarity with both themselves and others.
+Additionally, the familiarity may stem from overlapping training data used during pre-training stages, as many LLMs, including GPT-4, share similar datasets, influencing their output styles and content.
 4.2 Achieving high familiarity by exploring the target LLM generated responses
-Math Commensense Reasoning NLI
-Method Model Type GSM8K Math Algebra ECQA ESNLI
-Groundtruth Mistral 0.434 0.206 0.68 0.814
-Correct predicted sample + groundtruth 0.491 0.227 0.578 0.557
-correct preidcted sample only 0.449 0.257 0.586 0.698
-Style Transfered Groundtruth 0.51 19.6 0.696 0.845
-Table 4: Performance Comparison when not using GPT-4. For the style-transferred ground truth dataset, only 73.7% and 24% of the training data are used for training, as explained in Section5.2
+Method Model Type GSM8K Math Algebra ECQA
+Groundtruth Mistral 0.434 0.206 0.68
+Correct predicted sample + groundtruth 0.491 0.227 0.578
+Correct preidcted sample only 0.449 0.257 0.586
+Style Transfered Groundtruth 0.51 19.6 0.69
+Percentage of training data used in Style Transfered Groundtruth 73.7% 24% 91.2%
+Table 4: Performance Comparison when not using advanced LLMs. $n_{\text{train}}=1000$.
 In the experiments described above, we validated two hypotheses.
-However, all the studies mentioned use GPT-4 to generate responses, chosen for its advanced capabilities to directly produce correct answers and reliably follow diverse instructions.
+All the studies mentioned use advanced LLMs to generate responses, chosen for its advanced capabilities to directly produce correct answers and reliably follow diverse instructions.
 However, there is a concern regarding the "familiarity" factor: would this argument still hold without the involvement of an advanced LLM?
 If valid, the argument should apply to some extent even without using advanced LLMs.
-This section seeks to develop a method that enhances the “familiarity” of responses without requiring more advanced large language models (LLMs) than the target LLM.
-We included ESNLICamburu et al.
-(2018) as an additional dataset to better compare different data construction methods.
-The e-SNLI dataset enhances the SNLIBowman et al.
-(2015) dataset by adding human-annotated explanations to each instance.
-These annotations offer insights into the reasoning behind the labels of entailment, contradiction, or neutral, improving model interpretability and aiding in the development of more robust inference systems.
-Use the same model for train and groundtruth style transfer We devised a method to rewrite the ground truth data in a style that resembles language model-generated responses using the Mistral-7B-Instruct-V2 model.
+This section seeks to develop a method that enhances the “familiarity” of responses without requiring more advanced LLMs than the target LLM.
+Use the same model for train and groundtruth style transfer: We devised a method to rewrite the ground truth data in a style that resembles language model-generated responses using the Mistral-7B-Instruct-V2 model.
 The experiment results are shown in Table 4.
@@
 These two correct responses were then used as in-context examples to guide Mistral on how to rewrite ground truth data.
-Specifically, we used the two recorded predictions as example target labels and their corresponding questions as example questions, while the actual ground truth served as the example ground truth.
+Specifically, we used two recorded predictions as example target responses, their corresponding questions as example questions, and the related ground truth as the example ground truth.
 The in-context prompt informed Mistral that its task was to rewrite the ground truth in its own styles.
+For more details, please refer to Figure 6.
 Subsequently, we presented all questions and their corresponding ground truths from the training dataset to Mistral for style rewriting.
@@
 We noticed that even when provided with the ground truth, the Mistral model struggles to successfully rewrite answers for mathematical tasks.
-Therefore, we ran the rewritten math answers through our evaluation script and found that only 73.7% of the GSM8K answers and 24% of the math algebra answers passed the script.
-For math tasks, we only used the answers that passed the script for training.
-In the ECQA and ESNLI tasks, we allowed the Mistral model to rewrite the ground truth once and directly used all the transformed ground truth answers for training.
+Therefore, we ran the rewritten math answers through our evaluation script and found that only 73.7% of the GSM8K answers, 91.2% of the ECQA answers and 24% of the Math Algebra answers passed the script.
+We only used the answers that passed the script for training.
 This experiment was conducted solely with Mistral and not with the Llama model, due to our observations that Llama2-13b-chat lacks sufficient in-context following capabilities required for successful ground truth style transformation.
 We show a failure example in AppendixD.
-As a comparative experiment, we created the ’correct predicted samples only’ dataset.
-For each sample in this dataset, we had the Mistral model perform zero-shot generation 5 times, and using an evaluation filter, we randomly selected one correct prediction that passed the filter to be paired with the original input as the target label.
+As a comparative experiment, we created the ‘correct predicted samples only’ dataset.
+For each sample in this dataset, we had the Mistral model perform zero-shot generation 5 times, and using an evaluation filter, we randomly selected one correct prediction that passed the filter to be paired with the original input as a training sample.
 If the model failed to generate a correct answer in all 5 attempts, we did not include that data in the training dataset.
-The construction method for the ’correct predicted samples’ and ’correct predicted samples only’ datasets is similar.
+The construction method for the ‘correct predicted samples’ and ‘correct predicted samples only’ datasets is similar.
 The only difference is that if the sample fails to produce a correct answer, we include the ground truth in the dataset as the correct answer.
 As shown in Table 3, the Mistral model trained with style-transferred ground truth data outperformed the model trained with original ground truth data and other methods in all cases except for the Math Algebra dataset, where only 24% of the training data was used.
-To fairly compare the training effectiveness of the ’style transferred ground truth’ and the original ground truth, we recorded the IDs of the data used by the ’style transferred ground truth’.
+To fairly compare the training effectiveness of the ‘style transferred ground truth’ and the original ground truth, we recorded the IDs of the data used by the ‘style transferred ground truth’.
 We identified a corresponding set of 240 original ground truth data points for comparison.
-We found that training with only these 240 data points resulted in a 17.4% accuracy using the original ground truth, which is lower than the 19.6% achieved by the ’style transferred ground truth’.
-This demonstrates that data processed using the ’ground truth style transfer’ method yields better training outcomes than data labeled by humans.
-Interestingly, on the ECQA and ESNLI datasets, models trained with correct answers generated by Mistral in a zero-shot fashion performed worse than those trained directly with ground truth.
-By reading the ’correct predicted samples only’ dataset, we found that many instances classified as correct by the evaluation filter did not result from sound reasoning by the model.
-Although the final answers were technically correct, the underlying rationales were incorrect.
+We found that training with only these 240 data points resulted in a 17.4% accuracy using the original ground truth, which is lower than the 19.6% achieved by the ‘style transferred ground truth’.
+This demonstrates that data processed using the ‘ground truth style transfer’ method yields better training outcomes than data labeled by humans.
+Interestingly, on the ECQA dataset, models trained with correct answers generated by Mistral in a zero-shot fashion performed worse than those trained directly with ground truth.
+By reading the ‘correct predicted samples only’ dataset, we found that many instances classified as correct by the evaluation filter did not result from sound reasoning by the model.
 Therefore, the effectiveness of the self-training method depends significantly on whether the evaluation filter used to select correct predictions can truly identify accurate responses.
+Method Model Type GSM8K Math Algebra ECQA MBPP HumanEval Avg Perplexity Avg Token Length
+GPT-4 Answer Directly Mistra 0.597 0.302 0.722 0.354 0.365 3.81 164.642
+Minimum Change on Mistral Predictions 0.562 0.314 0.699 0.354 0.409 2.47 133.944
+Minimum Change on LlaMA Predictions 0.547 0.297 0.709 0.364 0.427 3.51 132.323
+Average Token Length for Mistral Initial Predictions 152.993
+GPT-4 Answer Directly Llama 0.428 0.150 0.656 0.2 0.158 3.58 167.469
+Minimum Change on LlaMA Predictions 0.433 0.166 0.649 0.2 0.213 2.75 132.323
+Minimum Change on Mistral Predictions 0.402 0.161 0.647 0.218 0.183 3.32 133.945
+Average Token Length for Llama Initial Predictions 165.793
+Table 5: Comparing the experimental results of GPT4 and minimum change. $n_{\text{train}}=1000$.
 5 A hybrid approach: creating responses using “minimal change” principal
@@
 We then instruct GPT-4 to copy the correct portions of the prediction and revise any erroneous parts of the response.
-When correcting errors, GPT-4 is required to make minimal changes to accurately adjust the answer  [^1] Please note that achieving such a process requires an advanced LLM, as the LLM needs to fully understand the solution and follows the “minimal change” principal.
+When correcting errors, GPT-4 is required to make minimal changes to accurately adjust the answer.
 Note that this construction process ensures that the target LLM is “familiar” with the response as the response is largely generated from itself.
@@
 This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained.
-To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ’Minimum Change Prompt and Method’ section of the Appendix C, which includes two to three examples of minimal changes along with explanations for each modification.
-Method Model Type GSM8K Math Algebra ECQA MBPP HumanEval Avg Perplexity
-Groundtruth Mistral 0.434 0.206 0.68 0.276 0.323 15.03
-GPT-4 Answer Directly 0.597 0.302 0.722 0.354 0.365 3.81
-Minimum Change on Mistral Predictions 0.562 0.314 0.699 0.354 0.409 2.47
-Minimum Change on LLaMA predictions 0.547 0.297 0.709 0.364 0.427 3.51
-Groundtruth Llama 0.364 0.126 0.631 0.2 0.146 6.18
-GPT-4 Answer Directly 0.428 0.150 0.656 0.2 0.158 3.58
-Minimum Change on LLaMA predictions 0.433 0.166 0.649 0.2 0.213 2.75
-Minimum Change on Mistral Predictions 0.402 0.161 0.647 0.218 0.183 3.32
-Table 5: Comparing the experimental results of GPT4, groundtruth and minimum change
-Results from our ’Minimum Change’ experiments are presented in Table 3.
-The data shows consistently lower perplexity compared to ground truth data across all datasets, indicating a high level of model familiarity with LLM-generated data.
-Furthermore, the performance across various setups was not only similar (with in-domain performance differences within 5% on most datasets) but also comparable to models trained with direct GPT-4 responses.
-Although the data was minimally modified based solely on the outputs from Mistral/LLaMA, theoretically, the logic and quality of the modified answers, as well as the level of detail, should be inferior to answers directly generated by GPT-4.
+To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ‘Minimum Change Prompt and Method’ section of the Appendix C.
+Results from our ‘Minimum Change’ experiments are presented in Table 5.
+The data consistently show lower perplexity compared to ground truth data across all datasets, indicating a high level of model familiarity with LLM-generated data.
+Furthermore, performance across various setups was not only similar (with in-domain performance differences within 5% on most datasets) but also comparable to models trained with direct GPT-4 responses.
+Although the data was minimally modified based solely on outputs from Mistral/Llama2, theoretically, the logic and quality of the modified answers, as well as the level of detail, should be inferior to answers directly generated by GPT-4.
+To verify whether the minimum change approach minimally added details, we measured the overall average token length.
+We first calculated the average token length for each dataset—GSM8K, Math Algebra, ECQA, MBPP, and HumanEval—and then averaged these results to obtain a composite mean.
+We found that the target responses from the minimum change were not only significantly shorter than those from ‘direct GPT-4 answers’ but also shorter than the initial predictions from Mistral/Llama2.
 Surprisingly, on the HumanEval and Math Algebra tasks, the performance of models trained with minimally changed data even surpassed those trained with data generated by GPT-4.
@@
 This underscores the effectiveness of training LLMs on data distributions familiar to them, demonstrating that minimal changes to maintain the original structure can be as effective as more detailed alterations.
-For more information about this experiment, please refer to the Appendix.
-B
+For more experiment details, see Appendix B.
 6 Related Works
-There are many studies related to using LLMs to generate training data for training models.
-Self-Training: Several works utilize a model’s own predictions for self-training.
-For instance, STAR Zelikman et al.
-(2022) and REST Gulcehre et al.
-(2023) generates a dataset through sample production from the LLMs, subsequently utilizing these samples to enhance the LLMs via training.
-RESTem Singh et al.
-(2023) enhances model performance by using initial predictions, filtering for accuracy, and retraining the model with correct predictions.
-This iterative process improves the model’s accuracy over multiple cycles.
-Data Augmentation: Numerous studies have investigated the use of synthetic data primarily for data augmentation purposes.
-In scenarios where sample sizes are insufficient, these studies generate additional training data to augment existing datasets (usually when the original training data is not adequate) Dai et al.
+Training on synthetic data has been widely employed to enhance model performance, particularly when training data is scarce.
+Various studies have focused on generating additional data to augment existing datasets Dai et al.
 (2023); Edwards et al.
@@
 (2023); Ubani et al.
-(2023); Agrawal et al.
-(2023); Piedboeuf and Langlais (2023), or for generating extensive training data when high-quality human-annotated instruction tuning data is difficult to obtain Kieser et al.
+(2023); Piedboeuf and Langlais (2023a); Agrawal et al.
+(2023); Piedboeuf and Langlais (2023b); Yu et al.
+(2024) or to create large-scale instruction-tuning datasets when high-quality human annotations are limited Kieser et al.
 (2023); Taori et al.
@@
 (2024).
-In cases where training data quality is low, many studies improve model training by providing more rationales Zhang et al.
-(2024); Kang et al.
-(2023); Li et al.
-(2022) or step-by-step reasoning Hsieh et al.
+In contrast, our work argues that for tasks requiring chain-of-thought reasoning, given the same question-answer pairs from the training dataset, the equivalent answers generated by LLMs are theoretically more effective than human-written ground truth due to familiarity.
+Previous works discuss various factors that improve training performance in reasoning, including adding complexity Xu et al.
+(2023), increasing diversity Luo et al.
+(2023), incorporating step-by-step reasoning Hsieh et al.
 (2023); Ho et al.
@@
 (2023); Fu et al.
-(2023) to small models.
-Our work distinguishes itself from existing studies by focusing on the role of familiarity in training.
-We explore the effectiveness of synthetic data from the perspective of familiarity, providing a theoretical foundation that substantiates the efficacy of past methods using synthetic data.
-Our finding offers a nuanced perspective that underscores the importance of dataset familiarity in enhancing model performance.
+(2023), adding details Zhang et al.
+(2024); Kang et al.
+(2023); Li et al.
+(2022), and ensuring correctness Trinh et al.
+(2024).
+We argue that familiarity is another important factor that can improve training outcomes, which has been overlooked by previous works.
+Several previous works share similarities with aspects of our research.
+For example, self-training methods like STAR Zelikman et al.
+(2022), REST Gulcehre et al.
+(2023), and RESTem Singh et al.
+(2023) generate datasets by producing samples from LLMs, filtering out mispredicted samples, and subsequently enhancing the models through further training on these filtered samples.
+The self-distillation method Yang et al.
+(2024) uses the target language model to rewrite the ground truth labels in its own words.
+While their self-distillation or rewriting processes are similar to the methods we employ in Section 4.2, they do not explore using models other than the target model.
+Our hypothesis suggests that data generated by any LLM is inherently familiar to the target LLM, indicating that more powerful models, such as GPT-4, could be used to generate rewritten labels instead of relying solely on the smaller target model being fine-tuned.
+Furthermore, these works do not identify the impact of the familiarity.
+Similar to our minimum-change approach, the works on learning from mistakes Singh et al.
+(2023) and process supervision Lightman et al.
+(2023); Luo et al.
+(2024) also involve steps of correcting a target LLM generated data.
+However, their overall process is different from our minimum-change.
+Moreover, they do not recognize the potential contribution of the familiarity in our approach and our work is complimentary to their discoveries.
+While they focus on learning from mistakes, our approach centers on maximizing familiarity.
+Perplexity has been used in previous studies primarily for data selection and filtering.
+Gonen et al.
+(2022) use perplexity to select prompts, demonstrating that lower perplexity prompts lead to better zero-shot QA performance on frozen models.
+Others have used perplexity to select pre-training data De la Rosa et al.
+(2022), detect AI-generated code Xu and Sheng (2024), and predict the synthetic generalization of language models Hu et al.
+(2020).
+Perplexity has also been employed to detect language model attacks Alon and Kamfonas (2023) and to select instruction tuning data Li et al.
+(2024); Mekala et al.
+(2024), where it reflects data difficulty.
+In contrast, we use perplexity to measure familiarity.
+Our focus is on how different outputs $y$, given the same input $x$, can affect training effectiveness due to variations in perplexity.
+We use perplexity to guide the construction of target responses that are more familiar to the model.
 7 Conclusion
-We systematically proved two hypotheses: LLMs are more familiar with responses generated by other LLMs when given similar inputs, and LLMs perform better on data they are familiar with.
-These hypotheses also explain why data annotated by LLMs yields better training outcomes than human-annotated data.
-The data generated by LLMs is more familiar to the trained LLMs, and training on unfamiliar data typically results in poorer outcomes.
-Based on this theory of familiarity, we have proposed several methods to generate data with high familiarity.
-Without using advanced LLMs, models can perform style transfers on ground truth to rewrite it in their own style.
-When using advanced LLMs for data processing, we can either use GPT-4 to read questions and generate answers in its own style, or employ our minimum change data generation method, which in some tasks, yields even better training results than direct responses from GPT-4.
+We conducted a series of experiments involving the rewriting of target responses to validate our two proposed hypotheses: 1) LLMs are more familiar with responses generated by themselves or other LLMs when given the same question; 2) LLMs exhibit improved performance when trained on data with which they are already familiar.
+These hypotheses help explain why data annotated by LLMs leads to superior training outcomes compared to human-annotated data on reasoning tasks, from a familiarity perspective.
+These findings suggest new directions for optimizing data annotation processes to enhance model performance.
 8 Limitations
-Perplexity can serve as a straightforward tool for measuring familiarity, but it is not very precise.
-It only provides relatively accurate predictions of performance decline when there are substantial differences in perplexity levels.
-We need a better tool to predict a model’s familiarity with a dataset and how that familiarity impacts the model’s learning.
-There may be additional forms of familiarity, such as linguistic familiarity and logical familiarity.
-Accurately identifying and measuring the impact of these forms of familiarity presents a challenge.
-Appendix A Appendix: Comparing varients of GPT-4 created training datasets
-GPT-4 Answer Directly: For mathematics and coding problems, we directly present the problems from the training dataset to GPT-4 for answers.
-For classification tasks, we provide GPT-4 with the gold-label (which does not include human-annotated rationales) and the input, and then use its responses.
-We combine these answers with the original questions to create the GPT-4 Answer Training Dataset.
-We purposefully do not supply GPT-4 with the solutions or rationales for the math, coding, or classification problems to prevent it from merely replicating the ground truth’s problem-solving and analytical processes.
+The effectiveness of the ground truth transformation method and the minimum change method can be limited by their accuracy.
+When an LLM is not powerful enough, it may not be able to successfully rewrite the ground truth in its own style while maintaining the correctness of the ground truth.
+Additionally, when performing minimal changes to a model’s initial prediction, we notice that GPT-4 often fails to correct the initial prediction accurately.
+Although GPT-4 may correct some errors, the minimally changed response can still be incorrect.
+Furthermore, minimal change is a challenging task.
+Achieving such a process requires an advanced LLM, as the LLM needs to fully understand the solution and follows the “minimal change” principal.
+Appendix A Comparing varients of GPT-4/Claude 3.5 created training datasets
+We instruct GPT-4/Claude 3.5 to generate target responses in various ways, and use these labels for training.
+The training results are summarized in Table 6 and Table 7, and we describe the different methods used to instruct GPT-4/Claude 3.5 to generate target responses.
+GPT-4/Claude 3.5 Answer Directly: For mathematics and coding problems, we directly present the problems from the training dataset to GPT-4/Claude 3.5 for answers.
+For classification tasks, we provide GPT-4/Claude 3.5 with the gold-label (which does not include human-annotated rationales) and the input questions, and then use its responses.
+We combine these answers with the original questions to create the GPT-4/Claude 3.5 Answer Training Dataset.
+We purposefully do not supply GPT-4/Claude 3.5 with the solutions or rationales for the math, coding, or classification problems to prevent it from merely replicating the ground truth’s problem-solving and analytical processes.
 If these processes were included, it could lead to a portion of GPT-4’s predictions not being generated in its own style, thus rendering our perplexity measurements—which assess how well an LLM handles predictions from other LLMs—less accurate.
-Step by step transformation on GT: For classification tasks, we provide GPT-4 with both the questions and the ground truth target labels along with human-annotated rationales.
+The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Response (No GT) response.
+"""We have the {question}
+1. We wish you to answer the question.
+2. You must answer the question (with inference process) directly without say anything else. Please not saying anything 'like sure I can help you with' or 'sure, i will not mention the gold label'
+3. You will inference first then put the Final Answer (NUMBER_HERE) at the end of the prediction like this
+INFERENCE HERE
+Final Answer: NUMBER_HERE"""
+Rewrite Ground Truth: For this method, we present GPT-4/Claude 3.5 with the ground truth data, including human-annotated rationales and problem-solving steps.
+The objective is to have GPT-4/Claude 3.5 rewrite the ground truth using its own language style.
+The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Response (Rewrite GT) response.
+"""Given the question: {question}
+and the groundtruth: {groundtruth}
+Please states the prediction in your own words. The groundtruth is 100% correct. You should not change the problem solving logic of the groundtruth. just restates it in your own words.
+1. You will pretend as you do not know the groundtruth, because we will use your prediction as target labels to train our model.
+2. (important format) You must generate the groundtruth directly. Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
+3. (important format) Please make sure the Final Answer: {gold_label} is placed at the end of the modified prediction."""
+Step-by-step: We specificly tell GPT-4/Claude 3.5 to solve the problem step by step.
+For mathematics and coding problems, we directly present the problems from the training dataset to GPT-4/Claude 3.5 for answers.
+For classification tasks, we provide GPT-4/Claude 3.5 with the gold-label (which does not include human-annotated rationales) and the input questions, and then use its step by step responses.
+We combine these answers with the original questions to create the GPT-4/Claude 3.5 Step-by-Step Response (No GT) Dataset.
+We purposefully do not supply GPT-4/Claude 3.5 with the solutions or rationales for the math, coding, or classification problems to prevent it from replicating the ground truth’s problem-solving and analytical processes.
+If these processes were included, it could lead to a portion of GPT-4/Claude 3.5’s predictions not being generated in its own style, thus rendering our perplexity measurements—which assess how well an LLM handles predictions from other LLMs—less accurate.
+The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Step-by-Step Response (No GT) response.
+We have the question and the groundtruth. Please reformat the groundtruth in step by step manner with details.
+Question: {question}
+Groundtruth: {groundtruth}
+1. We wish you to regenerate a new groundtruth. The new groundtruth solve the problem step by step. If you believe the groundtruth is not detail enough, you could add details.
+2. You will pretend as you do not know the groundtruth, because we will use your prediction as target labels to train our model.
+3.
+(important format) You must generate the groundtruth with the step by step inference process directly.
+Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
+4. (important format) You will inference first then put the Final Answer: {gold_label}
+at the end like this
+INFERENCE HERE
+Final Answer: {gold_label}
+Step-by-Step Transformation of GT: For classification tasks, we provide GPT-4/Claude 3.5 with both the questions and the ground truth target labels along with human-annotated rationales.
 For mathematical problems, we present the problems along with their answers, including the solutions process.
-We instruct GPT-4 to rewrite the human-annotated solutions or rationales in a step-by-step format, adding details where necessary to fill any gaps.
+We instruct GPT-4/Claude 3.5 to rewrite the human-annotated solutions or rationales in a step-by-step format, adding details where necessary to fill any gaps.
 If the existing content is already detailed enough, no additional details are added.
-We also instruct GPT-4 to preserve the original linguistic style as closely as possible by frequently incorporating language from the original ground truth.
-Detailed Step-by-Step Transformation of Ground Truth: This data construction method is very similar to the Step by Step method.
-The only difference is that we specifically instruct GPT-4 to add extra details when the original ground truth lacks sufficient information.
-Add details to GT: This approach involves enhancing the existing ground truth data by supplementing it with additional details to provide a richer context when necessary.
-We direct GPT-4 to enhance the content while maintaining the inherent language style of the original data.
-This method is applied selectively, only when the original information lacks sufficient clarity or depth to support effective model training and understanding.
-Rewrite GT in GPT-4 style words: For this method, we present GPT-4 with the ground truth data, including human-annotated rationales and problem-solving steps.
-The objective is to have GPT-4 rewrite the ground truth using its own language style.
-Datasets with explicit reasoning
-Math Commensense Reasoning NLI
-Method Training Dataset and Model Type GSM8K Math Algebra ECQA Perplexity Toekn Count
-Groundtruth GSM8K, Mistral 0.434 0.162 0.594 4.72 128.882
-GPT-4 Answer Directly 0.597 0.246 0.597 2.34 179.941
-Rewrite GT in GPT-4 style words 0.471 0.192 0.6 4.77 156.696
-Step by step 0.574 0.224 0.591 2.27 282.9
-Add details to GT 0.59 0.235 0.596 3.80 272.844
-Step by step transformation of GT 0.499 0.190 0.604 2.67 235.892
-Detailed Step-by-Step Transformation of Ground Truth 0.506 0.196 0.602 2.59 315.028
-Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.882
-GPT-4 Answer Directly 0.428 0.127 0.575 2.30 179.941
-Rewrite GT in GPT-4 style words 0.394 0.127 0.58 3.65 156.696
-Step by step 0.402 0.134 0.553 2.19 290.2
-Add details to GT 0.38 0.128 0.579
-Step by step transformation of GT 0.394 0.146 0.573 2.42 235.892
-Detailed Step-by-Step Transformation of Ground Truth 0.396 0.114 0.553 2.37 315.028
-Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.93 163.355
-GPT-4 Answer Directly 0.553 0.301 0.608 2.20 328.764
-Step by step 0.533 0.239 0.583 2.25 452
-Add details to GT 0.445 0.230 0.57 3.148 417.35
-Step by step transformation of GT 0.444 0.210 0.577 2.65 370.349
-Detailed Step-by-Step Transformation of Ground Truth 0.468 0.223 0.582 2.62 484.368
-Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.17 163.355
-GPT-4 Answer Directly 0.35 0.150 0.561 2.591 328.764
-Step by step 0.344 0.144 0.593 2.55 450.1
-Add details to GT 0.38 0.128 0.579 3.31 417.35
-Step by step transformation of GT 0.376 0.140 0.565 2.80 370.349
-Detailed Step-by-Step Transformation of Ground Truth 0.299 0.110 0.545 2.816 484.368
-Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.28 73.508
-GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.536
-Rewrite GT in GPT-4 style words 0.384 0.194 0.721 9.26 134.338
-Step by step 0.481 0.203 0.71 4.66 325.2
-Add details to GT 0.422 0.204 0.731 9.08 238.292
-Step by step transformation of GT 0.462 0.194 0.681 6.858 218.74
-Detailed Step-by-Step Transformation of Ground Truth 0.487 0.186 0.68 5.61 322.855
-Groundtruth ECQA, Llama 0.132 0.079 0.631 12.16 73.508
-GPT-4 Answer Directly 0.379 0.155 0.656 5.33 176.536
-Rewrite GT in GPT-4 style words 0.282 0.115 0.664 6.84 134.338
-Step by step 0.363 0.140 0.648 4.46 337.3
-Add details to GT 0.062 0.097 0.663 8.47 238.292
-Step by step transformation of GT 0.363 0.139 0.648 5.54 218.748
-Detailed Step-by-Step Transformation of Ground Truth 0.135 0.106 0.66 5.08 322.855
-Table 6: Appendix A: Additional experiments for Table2
-Appendix B Appendix: Minimum Change Vs GPT4 answer
-Datasets with explicit reasoning
-Math Commensense Reasoning NLI Code
-Method Training Dataset and Model Type GSM8K Math Algebra ECQA ESNLI HumanEval MBPP Perplexity Toekn Count
-Groundtruth GSM8K, Mistral 0.434 0.162 0.594 0.597 4.53 128.88
-GPT-4 Answer Directly 0.597 0.246 0.597 0.587 2.34 179.941
-Mistral Minimum Change Data 0.562 0.234 0.597 0.558 1.67 177.156
-llama minimum change data 0.547 0.242 0.621 3.19 166.152
-Groundtruth GSM8K, Llama 0.371 0.143 0.566 0.559 3.27 130.856
-GPT-4 Answer Directly 0.428 0.127 0.575 0.568 2.30 183.852
-Llama Minimum Change Data 0.433 0.140 0.602 0.602 2.22 166.15
-Mistral Minimum Change Data 0.402 0.141 0.576 2.249 177.156
-Groundtruth Math algebra, Mistral 0.264 0.206 0.554 0.574 5.83 163.355
-GPT-4 Answer Directly 0.553 0.301 0.608 0.555 2.19 328.764
-Mistral Minimum Change Data 0.536 0.313 0.622 0.558 1.85 297.609
-llama minimum change data 0.546 0.296 0.63 2.47 269.688
-Groundtruth Math algebra, Llama 0.36 0.126 0.575 0.552 5.14 162.417
-GPT-4 Answer Directly 0.35 0.150 0.561 0.564 2.59 326.573
-Llama Minimum Change Data 0.412 0.166 0.602 0.554 2.268 269.68
-Mistral Minimum Change Data 0.32 0.160 0.568 2.53 297.609
-Groundtruth ECQA, Mistral 0.258 0.1348 0.68 0.544 51.29 73.508
-GPT-4 Answer Directly 0.462 0.223 0.722 0.589 6.35 176.53
-Mistral Minimum Change Data 0.433 0.192 0.699 0.542 3.798 112.591
-LLama Minimum Change Data 0.483 0.212 0.709 4.89 128.695
-Groundtruth ECQA, Llama 0.132 0.079 0.631 0.431 11.95 76.118
-GPT-4 Answer Directly 0.379 0.155 0.656 0.479 5.33 183.182
-Llama Minimum Change Data 0.392 0.128 0.649 0.522 3.619 128.695
-Mistral Minimum Change Data 0.28 0.126 0.647 5.58 112.591
-Groundtruth ESNLI, Mistral 0.417 0.172 0.594 0.814 378.7 29.925
-GPT-4 Answer Directly 0.464 0.214 0.592 0.804 6.54 137.967
-Mistral Minimum Change Data 0.453 0.206 0.602 0.831 3.38 82.367
-Llama Minimum Change Data 0.432 0.203 0.63 0.836 4.15 97.078
-Groundtruth ESNLI, Llama 0.332 0.125 0.582 0.751 1201.70 29.963
-GPT-4 Answer Directly 0.386 0.123 0.564 0.776 5.28 143.739
-Llama Minimum Change Data 0.357 0.128 0.578 0.796 3.74 97.078
-Mistral Minimum Change Data 0.346 0.130 0.546 0.811 4.34 82.367
-Groundtruth HumanEval, Mistral 0.362 0.191 0.583 0.531 0.323 5.69 71.60
-GPT-4 Answer Directly 0.313 0.162 0.581 0.442 0.365 4.71 107.9
-Mistral Minimum Change Data 0.4055 0.202 0.5915 0.541 0.408 2.21 109.073
-Llama Minimum Change Data 0.4115 0.179 0.5835 0.5505 0.426 2.451 122.0
-Groundtruth HumanEval, Llama 0.0705 0.083 0.528 0.339 0.146 5.75 70.8
-GPT-4 Answer Directly 0.1245 0.105 0.553 0.473 0.158 4.33 107.14
-Llama Minimum Change Data 0.2445 0.389 0.5705 0.5175 0.213 2.21 122.0
-Mistral Minimum Change Data 0.131 0.103 0.528 0.4235 0.183 2.75 109.07
-Groundtruth MBPP, Mistral 0.392 0.175 0.527 0.481 0.276 7.78 71.60
-GPT-4 Answer Directly 0.399 0.186 0.568 0.553 0.354 3.419 786
-Mistral Minimum Change Data 0.405 0.179 0.601 0.48 0.354 2.817 63.86
-Llama Minimum Change Data 0.332 0.179 0.56 0.511 0.364 4.55 54.96
-Groundtruth MBPP, Llama 0.328 0.131 0.556 0.537 0.2 4.74 73.90
-GPT-4 Answer Directly 0.351 0.139 0.574 0.553 0.202 3.33 69.78
-Llama Minimum Change Data 0.34 0.130 0.569 0.539 0.204 3.44 54.96
-Mistral Minimum Change Data 0.333 0.123 0.563 0.544 0.218 3.45 63.86
-Table 7: Appendix B: More results for Table 5. Comparing the experimental results of GPT4, groundtruth and minimum change
-Appendix C Appendix: Minimum Change Prompt and Method
-Figure 3: Appendix C: Minimum Change Data Correction Examples
+We also instruct GPT-4/Claude 3.5 to preserve the original linguistic style as closely as possible by frequently incorporating language from the original ground truth.
+The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Step-by-Step Response (Rewrite GT) response.
+"""We have the question and the groundtruth.
+Please reformat the groundtruth in step by step manner.
+You must try your best to keep the original words and logic unchange.
+You are reformating the groundtruth into step by step solution.
+you could add new words, but you are trying to keep original words from groundtruth unchanged.
+Question: {question}
+Groundtruth: {groundtruth}
+1.
+We wish you to reformat a new groundtruth.
+The new groundtruth are reformated a new groundtruth which solve the problem step by step.
+you could add new words, but you are trying to not change the original words.
+2. You will pretend as you do not know the groundtruth, because we will use your prediction as target labels to train our model.
+3. When adding step by step inference process, you should imitate the language styles of the original groundtruth.
+4.
+(important format) You must generate the groundtruth with the step by step inference process directly.
+Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
+5. (important format) You will inference first then put the Final Answer: {gold_label}
+at the end like this
+INFERENCE HERE
+Final Answer: {gold_label}"""
+Detailed Step-by-Step Transformation of GT: This data construction method is very similar to the Step by Step method.
+The only difference is that we specifically instruct GPT-4/Claude 3.5 to add details when the original ground truth lacks sufficient information.
+The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Step-by-Step Detailed Response (Rewrite GT) response.
+"""We have the question and the groundtruth.
+Please reformat the groundtruth in step by step manner with details for each step.
+When adding detail, you must try your best to keep the original words and logic unchange.
+You are reformating the groundtruth with more details on each step instead of rewrite the groundtruth, but you are trying to keep original words from groundtruth unchanged.
+Question: {question}
+Groundtruth: {groundtruth}
+1. We wish you to reformat a new groundtruth. The new groundtruth are reformated a new groundtruth which solve the problem step by step with detail on each step.
+2. You will pretend as you do not know the groundtruth, because we will use your prediction as target labels to train our model.
+3. When adding step by step inference process, you should imitate the languaguage styles of the original groundtruth.
+4.
+(important format) You must generate the groundtruth with the detailed step by step inference process directly.
+Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
+5. (important format) You will inference first then put the Final Answer: {gold_label}
+at the end like this
+INFERENCE HERE
+Final Answer: {gold_label}"""
+GPT-4/Claude 3.5 Answer with Higher Perplexity: In this method, we first create a ‘paraphrased GPT-4/Claude 3.5 answer’ dataset.
+Specifically, we provide GPT-4 with the questions and the target responses from the ‘GPT-4/Claude 3.5 Response (No GT)’, instructing it to paraphrase only the target responses.
+We include the questions in the paraphrasing instruction prompt to ensure that the keywords in the paraphrased answers match those in the questions.
+This approach helps maintain content relevance.
+We then measure the perplexity of the ‘paraphrased GPT-4/Claude 3.5 response’ on the model that will be trained.
+We compare the perplexity of the target responses from the ‘GPT-4/Claude 3.5 Response (No GT)’ to the target responses from the ‘paraphrased GPT-4 response’.
+We select the one with the higher perplexity and collect them paired with the original question to form a new dataset, named ‘GPT-4/Claude 3.5 Answer with Higher Perplexity’.
+The following prompt instructs GPT-4/Claude 3.5 to create the paraphrased response.
+"""Given the question: {question}
+and the prediction: {gpt4_prediction}
+1. Please paraphrase the prediction. The prediction might be wrong sometimes, but you do not need to correct it. just paraphrase it.
+2.
+(important format) You must create the paraphrased prediction directly without say anything else.
+Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
+4. (important format) Please make sure the Final Answer: {gold_label_type} is placed at the end, which means you will not paraphrase the phrase 'Final Answer'."""
+Training Data Construction for Coding Tasks
+When constructing training datasets for coding tasks (such as HumanEval and MBPP), we follow a specialized pipeline.
+To build training datasets generated by models like GPT-4,Claude 3.5, Llama2, or Mistral, we first feed the problem to an LLM to generate an initial prediction.
+Next, we use GPT-4 to extract the corresponding code.
+These extracted codes are then assembled into target responses.
+The specific code extraction prompt is detailed below.
+f"""You need to extract the code from the previous prediction. The prediction answer the QUESTION.
+QUESTION: {modified_question}
+Previous prediction: {previous_prediction}
+Explaination of extract code:
+1.
+You have to extract code from the previous prediction.
+Your answer will be evaluated to see if it will pass the test, thus of course you need to extract the code and comments without including any other words.
+2. You only need to extract the final code for the fuction for the test case{item['test_list'][0]}. You do not care about the test cases.
+3. Please directly show me the extracted code with no other words. We will run the code to test if it is correct, thus do not extract other words.
+4. you suppose to extract the code from the previous prediction. You should never correct the code on your own even if it is wrong.
+Appendix B Minimum Change Vs GPT4 answer
+Table 8 summarizes the performance differences between GPT-4 and Minimum Change.
+According to the table, models trained on Minimum Change predictions perform comparably to those trained on GPT-4 generated predictions.
+The average length of the Minimum Change predictions is noticeably shorter than that of GPT-4 generated predictions for most tasks.
+Please note that Minimum Change is designed to maximize familiarity for the target model.
+For example, when testing on the HumanEval dataset, the cross-domain and in-domain performance of Llama’s Minimum Change data is noticeably higher than Mistral’s Minimum Change data when the target model is Llama.
+This is because Llama’s Minimum Change data focuses primarily on maximizing familiarity specifically for the Llama model.
+Appendix C Minimum Change Prompt and Method
+The minimum change pipeline is illustrated in Figure 3. The prompt used to instruct GPT-4 for creating the minimum change response is illustrated in Figure 4.
 Its construction process is as follows: For a given question, we first ask the target LLM, such as Llama2 or Mistral, to produce a response.
@@
 An illustration of such a modification is provided in Figure 3.
-We have detailed the prompt used to guide GPT-4 in the ’Minimum Change Prompt’ section of the Appendix.
-  • Training with Modified Predictions the minimum changed predictions are subsequently used as target labels for model training.
-Figure 4: Appendix C: Minimum Change Prompt Example
+We have detailed the prompt used to guide GPT-4 in the ‘Minimum Change Prompt’ section of the Appendix.
+  • Training with Modified Predictions the minimum changed predictions are subsequently used as target responses for model training.
 This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained.
-To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ’Minimum Change Prompt’ section of the Appendix, which includes two to three examples of minimal changes along with explanations for each modification.
-Appendix D Appendix: LLaMA groundtruth style transfer failure example
-The LLaMA-13b-chat model often fails to follow our in-context examples, leading to incorrect outputs when it is tasked with modifying ground truth data.
-Figure 5: Appendix D: LLaMA groundtruth style transfer failure example
+To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ‘Minimum Change Prompt’ section of the Appendix, which includes two to three examples of minimal changes along with explanations for each modification.
+Appendix D Llama2 groundtruth style transfer failure example
+The Llama2-13b-chat model often fails to follow our in-context examples, leading to incorrect outputs when it is tasked with modifying ground truth data. An example is illustrated in Figure 5.
+Appendix E Groundtruth Tranformation Prompt
+One of the example prompt is shown in Figure 6.
+Appendix F Detailed Evaluation Script
+Evaluation Script Details: To facilitate automatic evaluation across all models and datasets, we developed a specialized prompt format requiring models to conclude their predictions with ‘Final Answer: answer.’ For instance, the output should read ‘Final Answer: 100’ for an answer of 100.
+Our evaluation script extracts the text following the ‘Final Answer:’ substring.
+For the ECQA task, it verifies if the first text after ‘Final Answer:’ correctly matches one of the answer options.
+For mathematical problems including the GSM8K and MATH datasets, the script checks if the number after the substring ‘Final Answer:’ aligns with the ground truth number.
+We utilize the built-in testing script for the HumanEval dataset to ensure result accuracy and employ our custom scripts for all other tasks.
+By adapting the training data’s question and answer formats to this style, our models maintain consistent response formatting across different tasks.
+The evaluation scripts for HumanEval Chen et al.
+(2021) and MBPP Austin et al.
+(2021) require answers in a specific format.
+Models that have not been trained on this format struggle to follow it correctly, even when provided with in-context examples.
+Consequently, we do not conduct cross-domain evaluations on the MBPP and HumanEval.
+Appendix G More Setting Up Details
+We set the input token length to 512 and the output token length to 1024 for the GSM8K, MATH Algebra, and ECQA datasets.
+The reason for the extended output token length is that models trained on some of the GPT-4/Claud 3.5 response variants, such as ‘detailed step-by-step transformation of GT’ (as introduced in Appendix A), tend to generate lengthy predictions.
+We aim to ensure the output token length is sufficiently long.
+For instance, according to Table 6, the average token length for ‘detailed step-by-step transformation of GT’ in the Math Algebra dataset is 484.
+Clearly, a standard maximum output token length of 512 is not adequate for models trained on the ‘detailed step-by-step transformation of GT’ dataset.
+All of our code runs on a single A100 GPU.
+Appendix H Explanation of the Higher Perplexity of ECQA Human Responses
+The high-quality target responses are derived from the explanations in column "taskB" of the ECQA dataset (available at Hugging Face).
+The perplexity of the ECQA dataset is higher than that of math and coding datasets because textual reasoning tasks like ECQA involve more diverse response styles, leading to greater perplexity.
+Additionally, LLMs are extensively trained on math and coding tasks, making them more familiar with various response formats in those domains.
+Method Training Dataset and Model Type GSM8K Math Algebra ECQA Perplexity token length
+Zeroshot Mistral 0.413 0.185 0.504
+Zeroshot Llama2 0.35 0.125 0.597
+Ground truth GSM8K, Mistral 0.434 0.162 0.594 4.54 128.88
+GPT-4 Answer Directly 0.597 0.246 0.597 2.35 179.94
+GPT-4 Rewrite Ground Truth 0.471 0.192 0.6 4.77 156.696
+GPT-4 Step-by-step 0.574 0.224 0.591 2.27 282.9
+GPT-4 Step-by-Step Transformation of GT 0.499 0.190 0.6 2.67 235.892
+GPT-4 Detailed Step-by-Step Transformation of GT 0.506 0.196 0.602 2.59 315.028
+Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.88
+GPT-4 Answer Directly 0.428 0.128 0.575 2.30 179.941
+GPT-4 Rewrite Ground Truth 0.394 0.127 0.58 3.65 156.696
+GPT-4 Step-by-step 0.402 0.134 0.553 2.19 290.92
+GPT-4 Step-by-Step Transformation of GT 0.394 0.146 0.573 2.43 242.22
+GPT-4 Detailed Step-by-Step Transformation of GT 0.396 0.114 0.553 2.37 315.028
+Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.83 163.36
+GPT-4 Answer Directly 0.553 0.302 0.608 2.20 328.76
+GPT-4 Rewrite Ground Truth 0.545 0.310 0.57 3.51 267.11
+GPT-4 Step-by-step 0.533 0.239 0.583 2.250 451.99
+GPT-4 Step-by-Step Transformation of GT 0.444 0.210 0.577 2.65 370.35
+GPT-4 Detailed Step-by-Step Transformation of GT 0.468 0.223 0.582 2.62 484.368
+Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.15 163.36
+GPT-4 Answer Directly 0.35 0.150 0.561 2.59 328.76
+GPT-4 Rewrite Ground Truth 0.337 0.134 0.579 3.59 264.85
+GPT-4 Step-by-step 0.344 0.145 0.593 2.55 450.08
+GPT-4 Step-by-Step Transformation of GT 0.376 0.141 0.565 2.81 369.06
+GPT-4 Detailed Step-by-Step Transformation of GT 0.299 0.110 0.545 2.816 484.368
+Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.3 73.51
+GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.54
+GPT-4 Rewrite Ground Truth 0.384 0.194 0.721 9.269 134.338
+GPT-4 Step-by-step 0.481 0.203 0.71 4.66 325.15
+GPT-4 Step-by-Step Transformation of GT 0.462 0.194 0.681 6.89 218.748
+GPT-4 Detailed Step-by-Step Transformation of GT 0.487 0.186 0.68 5.61 322.855
+Groundtruth ECQA, Llama 0.132 0.0798 0.631 11.96 73.51
+GPT-4 Answer Directly 0.379 0.156 0.656 5.33 176.54
+GPT-4 Rewrite Ground Truth 0.282 0.116 0.664 6.84 134.338
+GPT-4 Step-by-step 0.363 0.140 0.648 4.46 337.30
+GPT-4 Step-by-Step Transformation of GT 0.312 0.108 0.628 5.55 226.99
+GPT-4 Detailed Step-by-Step Transformation of GT 0.135 0.106 0.66 5.08 322.855
+Table 6: Performance comparison of models trained on data constructed by different methods across various tasks.
+$n_{\text{train}}=1000$.
+Data points are labeled as low performance when their accuracy is more than 15% lower relative to the highest accuracy achieved in the same dataset using the same model.
+Method Training Dataset and Model Type GSM8K Math Algebra ECQA Perplexity token length
+Zeroshot Mistral 0.413 0.185 0.504
+Zeroshot Llama2 0.35 0.125 0.597
+Ground truth GSM8K, Mistral 0.434 0.162 0.594 4.54 128.88
+Claude Answer Directly 0.586 0.230 0.595 2.355 156.089
+Rewrite Ground Truth 0.614 0.213 0.536 2.900 159.561
+Claude Step-by-step 0.587 0.215 0.574 2.042 228.607
+Claude Step-by-Step Transformation of GT 0.49 0.205 0.59 2.309 208.132
+Claude Detailed Step-by-Step Transformation of GT 0.523 0.195 0.596 2.014 255.536
+Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.88
+Claude Answer Directly 0.433 0.110 0.548 2.037 159.291
+Claude Rewrite Ground Truth 0.361 0.130 0.557 2.30 163.0
+Claude Step-by-step 0.388 0.140 0.581 1.93 235.406
+Claude Step-by-Step Transformation of GT 0.399 0.141 0.568 2.08 213.658
+Claude Detailed Step-by-Step Transformation of GT 0.394 0.134 0.573 1.97 262.657
+Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.83 163.36
+Claude Answer Directly 0.554 0.277 0.606 2.236 245.346
+Claude Rewrite Ground Truth 0.5 0.242 0.595 2.801 219.599
+Claude Step-by-step 0.522 0.283 0.561 2.138 312.151
+Claude Step-by-Step Transformation of GT 0.446 0.203 0.579 2.319 285.7
+Claude Detailed Step-by-Step Transformation of GT 0.499 0.237 0.578 2.178 338.706
+Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.15 163.36
+Claude Answer Directly 0.317 0.137 0.54 2.33 244.672
+Claude Rewrite Ground Truth 0.272 0.120 0.554 2.70 218.456
+Claude Step-by-step 0.309 0.124 0.538 2.30 311.251
+Claude Step-by-Step Transformation of GT 0.235 0.105 0.51 2.424 284.719
+Claude Detailed Step-by-Step Transformation of GT 0.356 0.136 0.555 2.36 337.521
+Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.3 73.51
+Claude Answer Directly 0.457 0.213 0.714 4.02 198.55
+Claude Rewrite Ground Truth 0.415 0.201 0.689 6.21 124.47
+Claude Step-by-step 0.473 0.176 0.723 3.23 291.53
+Claude Step-by-Step Transformation of GT 0.47 0.195 0.694 4.69 215.99
+Claude Detailed Step-by-Step Transformation of GT 0.478 0.197 0.695 3.40 297.30
+Groundtruth ECQA, Llama 0.132 0.0798 0.631 11.96 73.51
+Claude Answer Directly 0.38 0.129 0.678 3.509 206.76
+Claude Rewrite Ground Truth 0.189 0.117 0.691 4.90 130.09
+Claude Step-by-step 0.334 0.126 0.644 3.36 302.60
+Claude Step-by-Step Transformation of GT 0.307 0.116 0.64 3.76 223.53
+Claude Detailed Step-by-Step Transformation of GT 0.168 0.096 0.637 3.33 308.76
+Table 7: Performance comparison of models trained on data constructed by different methods across various tasks.
+$n_{\text{train}}=1000$.
+Data points are labeled as low performance when their accuracy is more than 15% lower relative to the highest accuracy achieved in the same dataset using the same model.
+Math Commensense Reasoning Code
+Method Training Dataset and Model Type GSM8K Math Algebra ECQA HumanEval MBPP Perplexity Toekn Count
+Groundtruth GSM8K, Mistral 0.434 0.162 0.594 4.53 128.88
+GPT-4 Answer Directly 0.597 0.246 0.597 2.34 179.941
+Mistral Minimum Change Data 0.562 0.234 0.597 1.67 177.156
+llama minimum change data 0.547 0.242 0.621 3.19 166.152
+Mistral Initial Prediction Length 198.99
+Groundtruth GSM8K, Llama 0.371 0.143 0.566 3.27 130.856
+GPT-4 Answer Directly 0.428 0.127 0.575 2.30 183.852
+Llama Minimum Change Data 0.433 0.140 0.602 2.22 166.15
+Mistral Minimum Change Data 0.402 0.141 0.576 2.250 177.156
+Llama Initial Prediction Length 191.924
+Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.83 163.355
+GPT-4 Answer Directly 0.553 0.301 0.608 2.19 328.764
+Mistral Minimum Change Data 0.536 0.313 0.622 1.85 297.609
+llama minimum change data 0.546 0.296 0.63 2.47 269.688
+Mistral Initial Prediction Length 380.204
+Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.14 162.417
+GPT-4 Answer Directly 0.35 0.150 0.561 2.59 326.573
+Llama Minimum Change Data 0.412 0.166 0.602 2.268 269.68
+Mistral Minimum Change Data 0.32 0.160 0.568 2.53 297.609
+Llama Initial Prediction Length 297.576
+Groundtruth ECQA, Mistral 0.258 0.135 0.68 51.29 73.508
+GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.53
+Mistral Minimum Change Data 0.433 0.192 0.699 3.798 112.591
+LLama Minimum Change Data 0.483 0.212 0.709 4.89 128.695
+Mistral Initial Prediction Length 119.554
+Groundtruth ECQA, Llama 0.132 0.079 0.631 11.95 76.118
+GPT-4 Answer Directly 0.379 0.155 0.656 5.33 183.182
+Llama Minimum Change Data 0.392 0.128 0.649 3.619 128.695
+Mistral Minimum Change Data 0.28 0.126 0.647 5.58 112.591
+LLama Initial Prediction Length 149.945
+Groundtruth HumanEval, Mistral 0.362 0.191 0.583 0.323 5.69 71.60
+GPT-4 Answer Directly 0.313 0.162 0.581 0.365 4.71 107.9
+Mistral Minimum Change Data 0.406 0.202 0.592 0.408 2.21 109.07
+Llama Minimum Change Data 0.412 0.179 0.584 0.426 2.451 122.0
+Mistral Initial Prediction Length 118.80
+Groundtruth HumanEval, Llama 0.0705 0.083 0.528 0.146 5.75 70.8
+GPT-4 Answer Directly 0.125 0.105 0.553 0.158 4.33 107.14
+Llama Minimum Change Data 0.2445 0.389 0.5705 0.213 2.21 122.0
+Mistral Minimum Change Data 0.131 0.103 0.528 0.183 2.75 109.07
+Llama Initial Prediction Length 146.88
+Groundtruth MBPP, Mistral 0.392 0.175 0.527 0.276 7.78 71.60
+GPT-4 Answer Directly 0.399 0.186 0.568 0.354 3.419 78.6
+Mistral Minimum Change Data 0.405 0.179 0.601 0.354 2.817 63.86
+Llama Minimum Change Data 0.332 0.179 0.56 0.364 4.55 54.96
+Mistral Initial Prediction Length 66.22
+Groundtruth MBPP, Llama 0.328 0.131 0.556 0.2 4.74 73.90
+GPT-4 Answer Directly 0.351 0.139 0.574 0.202 3.33 69.78
+Llama Minimum Change Data 0.34 0.130 0.569 0.204 3.44 54.96
+Mistral Minimum Change Data 0.333 0.123 0.563 0.218 3.45 63.86
+Llama Initial Prediction Length 42.64
+Table 8: More results for Table 5.
+Comparing the experimental results of GPT4, groundtruth and minimum change.
+$n_{\text{train}}=1000$.
+The average token lengths for initial predictions by Mistral and Llama are shown in the table for GSM8K, Math Algebra, ECQA, MBPP(code only) and HumanEval(code only).
+Data points are labeled as low performance when their accuracy is more than 15% lower relative to the highest accuracy achieved in the same dataset using the same model.
+Figure 3: Minimum Change Data Correction Examples
+Figure 4: Minimum Change Prompt Example
+Figure 5: Llama2 groundtruth style transfer failure example
+Figure 6: Groundtruth Tranformation Prompt